### v2

# MOMENT: A Family of Open Time-series Foundation Models

Abstract
We introduce MOMENT, a family of open-source foundation models for general-purpose time series analysis. Pre-training large models on time series data is challenging due to (1) the absence of a large and cohesive public time series repository, and (2) diverse time series characteristics which make multi-dataset training onerous. Additionally, (3) experimental benchmarks to evaluate these models, especially in scenarios with limited resources, time, and supervision, are still in their nascent stages. To address these challenges, we compile a large and diverse collection of public time series, called the Time series Pile, and systematically tackle time series-specific challenges to unlock large-scale multi-dataset pre-training. Finally, we build on recent work to design a benchmark to evaluate time series foundation models on diverse tasks and datasets in limited supervision settings. Experiments on this benchmark demonstrate the effectiveness of our pre-trained models with minimal data and task-specific fine-tuning. Finally, we present several interesting empirical observations about large pre-trained time series models. Pre-trained models (AutonLab/MOMENT-1-large) and Time Series Pile (AutonLab/Timeseries-PILE) are available on https://huggingface.co/AutonLab.
Time series analysis, foundation models, pre-training, evaluation, forecasting, classification, anomaly detection, imputation
1 Introduction
Figure 1: MOMENT can solve multiple time series analysis tasks well (App. E).
Time series analysis is an important field encompassing a wide range of applications ranging from forecasting weather patterns (Schneider & Dickinson, 1974) or detecting irregular heartbeats using Electrocardiograms (Goswami et al., 2021), to identifying anomalous software deployments (Xu et al., 2018). Due to its significant practical value and the unique challenges that modeling time series data poses, time series analysis continues to receive substantial interest from academia and industry alike. However, modeling such data typically requires substantial domain expertise, time, and task-specific design.
Large pre-trained language (Touvron et al., 2023; Devlin et al., 2019; Chung et al., 2022), vision (Li et al., 2023a), and video (Day et al., 2023) models, typically perform well on a variety of tasks on data from diverse domains, with little or no supervision, and they can be specialized to perform well on specific tasks. We unlock these key capabilities for time series data and release the first family of open-source large pre-trained time series models, which we call MOMENT. The models in this family (1) serve as a building block for diverse time series analysis tasks (e.g., forecasting, classification, anomaly detection, and imputation, etc.), (2) are effective out-of-the-box, i.e., with no (or few) particular task-specific exemplars (enabling e.g., zero-shot forecasting, few-shot classification, etc.), and (3) are tunable using in-distribution and task-specific data to improve performance.
MOMENT is a family of high-capacity transformer models, pre-trained using a masked time series prediction task on large amounts of time series data drawn from diverse domains. Below we summarize our key contributions.
C1: Pre-training data. A key limiting factor for pre-training large time series models from scratch was the lack of a large cohesive public time series data repositories (Zhou et al., 2023; Gruver et al., 2023; Jin et al., 2023; Ekambaram et al., 2024; Cao et al., 2023). Therefore, we compiled The Time series Pile, a large collection of publicly available data from diverse domains, ranging from healthcare to engineering to finance. The Time Series Pile comprises of over 5 public time series databases, from several diverse domains for pre-training and evaluation (Tab. 11).
C2: Multi-dataset pre-training. Unlike text and images, which have largely consistent sampling rates and number of channels, time series frequently vary in their temporal resolution, number of channels [^1] Temporal resolution reflects sampling frequency of time series (e.g., hourly, daily); Channel is a single univariate time series in multivariate data (Ekambaram et al., 2024)., lengths, and amplitudes, and sometimes have missing values. As a result, large-scale mixed dataset pre-training is largely unexplored. Instead, most methods are trained on a single dataset, and transferred across multiple datasets, but with modest success (Wu et al., 2023; Oreshkin et al., 2021; Narwariya et al., 2020).
C3: Evaluation. Holistic benchmarks to evaluate time series foundation models on diverse datasets and tasks are in their nascent stages. Recent studies (Goswami et al., 2023b) have highlighted the importance of well-defined benchmarks and large-scale experimentation in order to accurately assess the impact and effectiveness of novel methodologies. To evaluate MOMENT, we build on the multi-task time series modeling benchmark first proposed by Wu et al. (2023) along multiple dimensions. For each of the 5 time series modeling tasks, namely, short- and long-horizon forecasting, classification, anomaly detection, and imputation we evaluate MOMENT against (1) both state-of-the-art deep learning as well as statistical baselines, on (2) more task-specific datasets, (3) using multiple evaluation metrics, (4) exclusively in limited supervision settings (e.g., zero-shot imputation, linear probing for forecasting, unsupervised representation learning for classification).
Finally, we explore various properties of these pre-trained time series models. In particular, we study whether MOMENT is aware of intuitive time series characteristics such as frequency and trend, and the impact of initialization, model size scaling, and cross-modal transfer.
2 Related Work
Transformers and patching for time series modeling. There is a growing body of work utilizing transformers for various time series analysis tasks (Wen et al., 2023). One issue with applying transformers to time series data is the complexity of the self-attention mechanism, which grows quadratically with the size of input tokens (or length of time series) (Li et al., 2019). Nie et al. (2023) demonstrated that treating time series sub-sequences (or patches) as tokens instead of individual time points is a simple, efficient, and effective mechanism for learning useful representations for forecasting. Drawing inspiration from prior work, we build on top of the transformer architecture which takes disjoint time series sub-sequences (or patches) as input.
Masked Representation Learning. Masked pre-training is a widely-used self-supervised learning task where a model learns to accurately reconstruct masked portions of its input. Masked language (Devlin et al., 2019; Raffel et al., 2020) and image modeling (Xie et al., 2022; Li et al., 2023b) have been successfully utilized to learn models from vast quantities of unlabeled data, which can generalize to a variety of downstream tasks.
For time series data, prior work has primarily focused on contrastive representation learning (Yue et al., 2022; Eldele et al., 2021; Franceschi et al., 2019). However, contrastive learning relies on data augmentation, which is both subjective and data-dependent. In contrast, some studies mask portions of time series using zeros and learn a model to reconstruct them (Nie et al., 2023; Zerveas et al., 2021; Dong et al., 2023; Li et al., 2023c).
Representation learning via masking is well-suited to all the downstream tasks we care about, especially forecasting and imputation, as they are instances of the masked reconstruction problem. Due to its simplicity and success in vision and language domains, we use the masked prediction task to pre-train our model, using a special embedding (see [MASK] in Fig. 3) to mask time series patches instead of zeros.
Cross-modal transfer learning using language models. Lu et al. (2022) had first shown that transformers pre-trained on text data (LLMs) can effectively solve sequence modeling tasks in other modalities. Subsequently, Shen et al. (2023) introduced ORCA, a general cross-modal fine-tuning framework that extends the applicability of a single large-scale pretrained model to diverse modalities by adapting to a target task via an align-then-refine workflow. Given the target input, ORCA first learns an embedding network that aligns the embedded feature distribution with the pretraining modality, then the pretrained model is fine-tuned on the embedded data, exploiting the knowledge shared across modalities. Some recent studies have leveraged this inherent ability of language pre-trained transformers to “reprogram” LLMs for time series analysis using parameter efficient fine-tuning and suitable tokenization strategies (Zhou et al., 2023; Gruver et al., 2023; Jin et al., 2023; Cao et al., 2023; Ekambaram et al., 2024). However, some of these models (Jin et al., 2023; Gruver et al., 2023) with billions of parameters demand significant memory and computational resources to perform well. We complement this line of research with three empirical observations (Sec 4.3): we show that (1) transformers trained on time series can also model sequences across modalities, (2) during pre-training, randomly initializing weights lead to lower pre-training loss, than initializing with language modeling weights, and (3) models pre-trained on time series outperform LLM-based models such as (Zhou et al., 2023; Jin et al., 2023) on many tasks and datasets.
Unanswered Questions. To the best of our knowledge, two questions remain largely unanswered in prior work on time series modeling. First, all existing time series models are (pre-)trained and fine-tuned on individual datasets (Nie et al., 2023; Yue et al., 2022; Wu et al., 2023; Zhou et al., 2023), and the benefits (or drawbacks) of large-scale multi-dataset pre-training remains unexplored (Wen et al., 2023). Second, there is very limited work on time series modeling in limited supervision settings, such as zero-shot forecasting (Oreshkin et al., 2021), or few-shot classification (Narwariya et al., 2020). In our work, we consider both these questions and show that pre-training a model of sufficient capacity on a large corpus of unlabeled time series data can in fact enable it to provide reasonably accurate predictions in limited-supervision settings.
3 Methodology
We first collect a large number of public time series data into the Time Series Pile and then use it to pre-train a transformer model on the masked time series prediction task. We discuss each of these steps in the following sections.
3.1 The Time Series Pile
Figure 2: Time Series Pile data splits. To avoid data contamination, we carefully partition all datasets into disjoint train, validation, and test splits. We adhere to the predefined splits provided by the creators of each dataset. In cases where such splits are unavailable, we randomly sample 60% of the data for training, 10% for validation, and 30% for testing. We only use the training splits of all datasets for pre-training.
Unlike natural language processing and computer vision, where large-scale datasets such as The Pile (Gao et al., 2020), and ImageNet-1K (Russakovsky et al., 2015) are easily available for pre-training, public time series datasets are much smaller, scattered, and largely task-specific (Ma et al., 2023; Zhou et al., 2023; Gruver et al., 2023). To bridge this gap, we collate multiple time series from 4 task-specific, widely-used public repositories resulting in a large number of time series spanning diverse domains, and time series characteristics such as lengths, amplitudes, and temporal resolutions. We call this collection the Time Series Pile.
Informer long-horizon forecasting datasets (Zhou et al., 2021) is a collection of 9 datasets that are widely used to evaluate long-horizon forecasting performance (Wu et al., 2023; Nie et al., 2023; Challu et al., 2023): 2 hourly and minutely subsets of the Electricity Transformer Temperature (ETT(Zhou et al., 2021), Electricity (Trindade, 2015), Traffic (California Department of Transportation, 2024), Weather (Max Planck Institute for Biogeochemistry, 2024), Influenza-like Illness (ILI(Centers for Disease Control and Prevention, 2024), and Exchange-rate (Lai et al., 2018).
Monash time series forecasting archive (Godahewa et al., 2021) is a collection of 58 publicly available short-horizon forecasting datasets with a total of over 100K time series, spanning a variety of domains and temporal resolutions.
UCR/UEA classification archive (Dau et al., 2018) comprises of 159 time series datasets which are frequently used to benchmark classification algorithms (Ismail Fawaz et al., 2019). These datasets belonging to seven different categories (Image Outline, Sensor Readings, Motion Capture, Spectrographs, ECG, Electric Devices, and Simulated Data), vary substantially in terms of the number of classes and the size of the training set.
TSB-UAD anomaly benchmark (Paparrizos et al., 2022b) is a recent collection of 1980 univariate time series with labeled anomalies from 18 anomaly detection datasets proposed over the past decade. This collection includes both synthetic and real-world time series originating from a wide range of sources such as the human body, spaceships, environment, and web serves.
Minimizing data contamination using careful train-test splitting. We carefully split each dataset into disjoint training, validation, and test splits, based on splits specified by data creators. When these splits are not available, we randomly sample 60% of the data for training, 10% for validation, and 30% for testing. Long-horizon forecasting and anomaly detection datasets are typically long time series, which are split horizontally as shown in Fig. 2. Conversely, short-horizon forecasting and classification datasets often contain multiple short time series. For these datasets, a complete time series is either training, validation, or testing. We use the same random seed, set to 13, throughout our experiments, from pre-training to downstream evaluation, thus ensuring that MOMENT only observes the training splits of datasets during pre-training.
3.2 Model Architecture
Figure 3: Overview of MOMENT. A time series is broken into disjoint fixed-length sub-sequences called patches, and each patch is mapped into a $D$-dimensional patch embedding. During pre-training, we mask patches uniformly at random by replacing their patch embeddings using a special mask embedding [MASK]. The goal of pre-training is to learn patch embeddings which can be used to reconstruct the input time series using a light-weight reconstruction head.
MOMENT receives a univariate time series $\mathcal{T}\in\mathbb{R}^{1\times T}$, and a mask $M=\{0,1\}^{1\times T}$ of length $T$. 0 and 1 denote unobserved and observed time-stamps, respectively. Reversible instance normalization (Kim et al., 2022) is applied to the observed time series before breaking it into $N$ disjoint patches of length $P$. Each patch is then mapped to a $D$-dimensional embedding, using a trainable linear projection if all time steps are observed, and a designated learnable mask embedding [MASK] $\in\mathbb{R}^{1\times D}$, otherwise. These $N$ patch embeddings serve as input to the transformer model which retains their shape (1 ×$\times D$) throughout its operations. Each transformed patch embedding is then used to reconstruct both masked and unmasked time series patches, using a lightweight prediction head. The goal of the prediction head is to map the transformed patch embeddings to the desired output dimensions. Since this particular prediction head enables time series reconstruction, we call it the reconstruction head. Fig. 3 shows an overview of our model.
Our transformer encoder retains the modifications proposed by Raffel et al. (2020) to the original Transformer (Vaswani et al., 2017). Specifically, we remove the additive bias from the Layer Norm (Ba et al., 2016), and place it before the residual skip connections (He et al., 2016), and use the relation positional embedding scheme (Shaw et al., 2018). Below we summarize the intuition behind some of our key design decisions.
Handling varying time series characteristics. Time series vary in length, number of channels, amplitudes, and temporal resolutions. We address variable length by restricting MOMENT’s input to a univariate time series of a fixed length $T=$ 512. As is common practice, we sub-sample longer time series, and pad shorter ones with zeros on the left [^2] We found a large majority of classification datasets to have time series shorter than 512. Besides, a look-back window of length 512 was found to be sufficient for accurate long-horizon forecasting (Nie et al., 2023).. Moreover, segmenting time series into patches quadratically reduces MOMENT’s memory footprint and computational complexity, and linearly increases the length of time series it can take as input. We handle multi-variate time series by independently operating on each channel along the batch dimension. Like recent studies (Zhou et al., 2023; Nie et al., 2023), we found that modeling each channel independently is an effective strategy for modeling multivariate time series. Finally, re-scaling and centering time series using reversible instance normalization enables MOMENT to model time series with significantly different temporal distributions (Kim et al., 2022). We did not explicitly model the temporal resolution of time series, since this information is often unavailable outside of time series forecasting datasets.
Intentionally simple encoder. Closely following the design of transformers in the language domain allows us to leverage their scalable and efficient implementations (e.g., gradient checkpointing, mixed precision training).
Light-weight prediction head. We use a lightweight prediction head instead of a decoder of the same size as the encoder, to enable the necessary architectural modifications for task-specific fine-tuning of a limited number of trainable parameters while keeping the bulk of parameters and the high-level features learned by the encoder intact.
Tasks Supervision Datasets Metrics Baselines Experimental Setting
Long-horizon
Forecasting
Linear Probing
ETT-h1/h2/m1/m2,
Electricity, Traffic,
Weather, Exchange, ILI
MSE, MAE
Time-LLM, GPT4TS,
TimesNet, PatchTST, FEDFormer,
DLinear, N-BEATS,
Stationary, LightTS
Look-back window $L=512$,
[…]
Short-horizon
Forecasting
Zero-shot
M3 and M4
competition
datasets (subset)
sMAPE [^3] The definitions of sMAPE were different in the M3 and M4 competitions. In our experiments, we used the same definition as the M4 competition.
GPT4TS, TimesNet, N-BEATS,
AutoARIMA, AutoTheta, AutoETS,
Seasonal Naive, Naive, Random Walk
Statistical methods fit on individual time series.
Deep learning methods are trained on a source dataset
& evaluated on a target dataset of the same temporal resolution.
Classification
Unsupervised
representation
learning
UCR Classification
Archive (subset)
Accuracy
GPT4TS, TimesNet,
TS2Vec, T-Loss, TNC, TS-TCC, TST,
CNN, Encoder, FCN, MCNN,
MLP, ResNet, t-LeNet, TWIESN
DTW
All models except MOMENT were trained on each
individual dataset. Quality of unsupervised representations
measured using the accuracy of a SVM trained on them.
Anomaly
Detection
Linear probing,
Zero-shot
UCR Anomaly
Archive (subset)
Adjusted Best F1
VUS-ROC
GPT4TS, TimesNet,
Anomaly Transformer, DGHL,
Anomaly Nearest Neighbor
Reconstruction-based anomaly detection with window size =512=512= 512
MSE between observed and predicted time series
is used as the anomaly criterion
Imputation
Linear probing,
Zero-shot
ETT-h1/h2/m1/m2,
Electricity, Weather
MSE, MAE
GPT4TS, TimesNet,
Linear, Naive, Cubic Spline, Nearest Neighbors
Randomly mask contiguous sub-sequences of
length 8
[…]
Table 1: Experimental benchmark. We evaluate MOMENT on 5 time series analysis tasks with an emphasis on limited memory, compute, and supervision settings.
3.3 Pre-training using Masked Time series Modeling
We pre-train MOMENT using the masked time series modeling task. Fig. 3 presents an overview of our pre-training procedure. During training, we first mask a small number of patches uniformly at random by replacing their patch embeddings with a learnable mask embedding [MASK]. The corrupted time series patches are then fed into the transformer encoder to learn patch representations, which are used to reconstruct the original time series using a lightweight reconstruction head. The pre-training objective is to minimize the masked reconstruction error i.e. the Mean Squared Error between the ground truth and the prediction, averaged over the masked patches.
Pre-training Setup. We pre-train three different sizes of MOMENT, roughly corresponding to the sizes of encoders in T5-Small, Base, and Large. Specifically, the Base (Small, $\overline{\mbox{{Large}}}$) model uses a 12 (6, $\overline{\mbox{24}}$) layer Transform with hidden dimensions of size $D=$ 768 (512, $\overline{\mbox{1024}}$), 12 (8, $\overline{\mbox{16}}$) attention heads, and feed-forward networks of size 3072 (2048, $\overline{\mbox{4096}}$), resulting in approximately 125 (40, $\overline{\mbox{385}}$) million parameters. All weights are randomly initialized before pre-training. All models take an input time series of length $T=$ 512, breaking it into $N=$ 64 disjoint patches of length $P=$ 8. We mask 30% of the patches uniformly at random during pre-training.
We use the Adam optimizer with weight decay (Loshchilov & Hutter, 2019) with $\lambda=$ 0.05, $\beta_{1}=$ 0.9, $\beta_{2}=$ 0.999. We clip the gradient at 5.0, train models using a batch size of 2048, and use cosine learning rate schedule with initial and final learning rates of $1e^{-4}$ and $1e^{-5}$, respectively. We use gradient checkpointing (Radford et al., 2021) to improve training throughput and save memory, and train all models in a mixed precision setting, using float-32 for numerically unstable operations, e.g. layer normalization, and bfloat-16444https://cloud.google.com/tpu/docs/bfloat16, otherwise. We train all models for 2 epochs.
3.4 Fine-tuning on Downstream Tasks
MOMENT can be seamlessly used for multiple time series analysis tasks. In this work, we consider 5 practical time series analysis tasks as examples, namely: long- and short-horizon forecasting, classification, anomaly detection, and imputation. For forecasting tasks with horizon $H$, we replace the reconstruction head with a forecasting head, which first flattens all the $N$ $D$-dimensional patch embeddings into a $N\times D$ dimensional vector, and then projects it into a $H$-dimensional time series via a linear projection layer. For all other tasks, we retain the reconstruction head. We provide detailed descriptions of each task and MOMENT’s configuration in App. E.
Fine-tuning settings. MOMENT can either be fine-tuned end-to-end, or linear probed ($\mathtt{{MOMENT_{LP}}}$) by freezing all parameters except for those in the reconstruction or forecasting head. Additionally, for some tasks such as anomaly detection, unsupervised representation learning and imputation, MOMENT can also be used in a zero-shot ($\mathtt{{MOMENT_{0}}}$) setting by retaining its reconstruction head.
4 Experimental Setup and Results
Methods $\mathtt{MOMENT_{LP}}$ Time-LLM GPT4TS PatchTST DLinear TimesNet FEDFormer Stationary LightTS N-BEATS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.154 0.209 - - 0.162 0.212 0.149 0.198 0.176 0.237 0.172 0.220 0.217 0.296 0.173 0.223 0.182 0.242 0.152 0.210
192 0.197 0.248 - - 0.204 0.248 0.194 0.241 0.220 0.282 0.219 0.261 0.276 0.336 0.245 0.285 0.227 0.287 0.199 0.260
336 0.246 0.285 - - 0.254 0.286 0.245 0.282 0.265 0.319 0.280 0.306 0.339 0.380 0.321 0.338 0.282 0.334 0.258 0.311
720 0.315 0.336 - - 0.326 0.337 0.314 0.334 0.333 0.362 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.386 0.331 0.359
ETTh1 96 0.387 0.410 0.408 0.429 0.376 0.397 0.370 0.399 0.375 0.399 0.384 0.402 0.376 0.419 0.513 0.491 0.424 0.432 0.399 0.428
192 0.410 0.426 - - 0.416 0.418 0.413 0.421 0.405 0.416 0.436 0.429 0.420 0.448 0.534 0.504 0.475 0.462 0.451 0.464
336 0.422 0.437 - - 0.442 0.433 0.422 0.436 0.439 0.443 0.491 0.469 0.459 0.465 0.588 0.535 0.518 0.488 0.498 0.500
720 0.454 0.472 0.523 0.514 0.477 0.456 0.447 0.466 0.472 0.490 0.521 0.500 0.506 0.507 0.643 0.616 0.547 0.533 0.608 0.573
ETTh2 96 0.288 0.345 0.285 0.348 0.285 0.342 0.274 0.336 0.289 0.353 0.340 0.374 0.358 0.397 0.476 0.458 0.397 0.437 0.327 0.387
192 0.349 0.386 - - 0.354 0.389 0.339 0.379 0.383 0.418 0.402 0.414 0.429 0.439 0.512 0.493 0.520 0.504 0.400 0.435
336 0.369 0.408 - - 0.373 0.407 0.329 0.380 0.448 0.465 0.452 0.452 0.496 0.487 0.552 0.551 0.626 0.559 0.747 0.599
720 0.403 0.439 0.399 0.435 0.406 0.441 0.379 0.422 0.605 0.551 0.462 0.468 0.463 0.474 0.562 0.560 0.863 0.672 1.454 0.847
ETTm1 96 0.293 0.349 0.384 0.403 0.292 0.346 0.290 0.342 0.299 0.343 0.338 0.375 0.379 0.419 0.386 0.398 0.374 0.400 0.318 0.367
192 0.326 0.368 - - 0.332 0.372 0.332 0.369 0.335 0.365 0.374 0.387 0.426 0.441 0.459 0.444 0.400 0.407 0.355 0.391
336 0.352 0.384 - - 0.366 0.394 0.366 0.392 0.369 0.386 0.410 0.411 0.445 0.459 0.495 0.464 0.438 0.438 0.401 0.419
720 0.405 0.416 0.437 0.429 0.417 0.421 0.416 0.420 0.425 0.421 0.478 0.450 0.543 0.490 0.585 0.516 0.527 0.502 0.448 0.448
ETTm2 96 0.170 0.260 0.181 0.269 0.173 0.262 0.165 0.255 0.167 0.269 0.187 0.267 0.203 0.287 0.192 0.274 0.209 0.308 0.197 0.271
192 0.227 0.297 - - 0.229 0.301 0.220 0.292 0.224 0.303 0.249 0.309 0.269 0.328 0.280 0.339 0.311 0.382 0.285 0.328
336 0.275 0.328 - - 0.286 0.341 0.274 0.329 0.281 0.342 0.321 0.351 0.325 0.366 0.334 0.361 0.442 0.466 0.338 0.366
720 0.363 0.387 0.366 0.388 0.378 0.401 0.362 0.385 0.397 0.421 0.408 0.403 0.421 0.415 0.417 0.413 0.675 0.587 0.395 0.419
ILI 24 2.728 1.114 3.025 1.195 2.063 0.881 1.319 0.754 2.215 1.081 2.317 0.934 3.228 1.260 2.294 0.945 8.313 2.144 4.539 1.528
36 2.669 1.092 - - 1.868 0.892 1.430 0.834 1.963 0.963 1.972 0.920 2.679 1.080 1.825 0.848 6.631 1.902 4.628 1.534
48 2.728 1.098 - - 1.790 0.884 1.553 0.815 2.130 1.024 2.238 0.940 2.622 1.078 2.010 0.900 7.299 1.982 4.957 1.585
60 2.883 1.126 3.245 1.221 1.979 0.957 1.470 0.788 2.368 1.096 2.027 0.928 2.857 1.157 2.178 0.963 7.283 1.985 5.429 1.661
ECL 96 0.136 0.233 - - 0.139 0.238 0.129 0.222 0.140 0.237 0.168 0.272 0.193 0.308 0.169 0.273 0.207 0.307 0.131 0.228
192 0.152 0.247 - - 0.153 0.251 0.157 0.240 0.153 0.249 0.184 0.289 0.201 0.315 0.182 0.286 0.213 0.316 0.153 0.248
336 0.167 0.264 - - 0.169 0.266 0.163 0.259 0.169 0.267 0.198 0.300 0.214 0.329 0.200 0.304 0.230 0.333 0.170 0.267
720 0.205 0.295 - - 0.206 0.297 0.197 0.290 0.203 0.301 0.220 0.320 0.246 0.355 0.222 0.321 0.265 0.360 0.208 0.298
Traffic 96 0.391 0.282 - - 0.388 0.282 0.360 0.249 0.410 0.282 0.593 0.321 0.587 0.366 0.612 0.338 0.615 0.391 0.375 0.259
192 0.404 0.287 - - 0.407 0.290 0.379 0.256 0.423 0.287 0.617 0.336 0.604 0.373 0.613 0.340 0.601 0.382 0.403 0.274
336 0.414 0.292 - - 0.412 0.294 0.392 0.264 0.436 0.296 0.629 0.336 0.621 0.383 0.618 0.328 0.613 0.386 0.426 0.285
720 0.450 0.310 - - 0.450 0.312 0.432 0.286 0.466 0.315 0.640 0.350 0.626 0.382 0.653 0.355 0.658 0.407 0.508 0.335
Table 2: Long-term forecasting performance measured using Mean Squared Error (MSE) and Mean Absolute Error (MAE). PatchTST performs the best across most settings, closely followed by MOMENT. Complete results in Tab. 18.
$\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet N-BEATS ARIMA Theta ETS
Seasonal
Naive
Naive
Random
Walk
Datasets M4 FR M4 FR M4 FR M4 FR
M3 Yearly 16.74 16.97 18.39 17.40 27.48 16.21 16.82 15.92 17.90 16.70 16.47 17.54 17.54 16.77
Quarterly 10.09 10.62 10.18 10.29 14.41 12.68 11.26 11.30 10.18 9.19 8.99 11.02 11.45 11.72
Monthly 16.04 16.90 15.21 16.37 15.58 16.23 15.63 16.37 15.95 14.96 14.41 17.74 18.53 19.19
M4 Yearly - 14.84 - 14.80 - 14.40 - 14.18 16.19 14.04 14.06 16.33 16.33 14.22
Quarterly - 12.02 - 11.77 - 13.21 - 12.25 10.86 10.21 10.24 12.55 11.65 11.46
Monthly - 15.80 - 15.36 - 15.67 - 15.24 13.68 13.19 13.58 16.00 15.24 15.48
Table 3: Zero-shot short-horizon forecasting performance on a subset of the M3 and M4 datasets measured using sMAPE. Statistical methods outperformed their deeper counterparts. However, on some datasets (in bold), MOMENT, GPT4TS and N-BEATS achieved a lower sMAPE than ARIMA.
We extend the experimental benchmark introduced by Wu et al. (2023) across various dimensions. Below, we outline the design choices of our benchmark and highlight its key distinctions from TimesNet [^5] In this section, we use TimesNet to refer to the benchmark proposed by Wu et al. (2023) instead of their model..
Time series modeling with limited supervision. Our benchmark comprises of 5 major time series modeling tasks of significant practical value, namely long- and short-horizon forecasting, imputation, classification, and anomaly detection, as outlined in Tab. 1. In contrast to TimesNet, we exclusively consider scenarios characterized by limited compute and supervision resources. These scenarios mimic practical situations where training (or fine-tuning) a deep neural network is infeasible due to resource limitations or insufficiently characterized data. Accordingly, we assess MOMENT in zero-shot settings whenever feasible and through linear probing for a few epochs otherwise.
For classification, we consider the unsupervised representation learning problem, where the goal is to learn representations of time series that are useful for downstream classification, without access to labeled data. As is common in prior work (Yue et al., 2022; Franceschi et al., 2019), the quality of representations is measured using the accuracy of a Support Vector Machine trained on them (App. E.2). For short-horizon forecasting, we consider the zero-shot setting introduced by Oreshkin et al. (2021). In particular, we fine-tune MOMENT on a source dataset using a forecasting head, and evaluate its performance on a target dataset without any fine-tuning (App E.1.2, Tab. 21).
Datasets. We use the same datasets as TimesNet for forecasting and imputation. However, for classification and anomaly detection, we conduct experiments on larger and systematically chosen subset of datasets from the UCR classification archive (Dau et al., 2018) and UCR anomaly archive (Wu & Keogh, 2023). Specifically, we run classification experiments on all 91 time series datasets with each time series shorter than 512 time steps (Tab.23). For anomaly detection, while choosing the subset of time series, we prioritized coverage over different domains and data sources represented in the UCR anomaly archive (Tab. 22). We also note that the UCR anomaly archive was proposed as an improvement over pre-existing anomaly detection datasets such as the SMD (Su et al., 2019), and SMAP (Hundman et al., 2018), many of which are also used in TimesNet. Our proposed experimental setup is summarized in Tab. 1 and detailed in App. E.
Metrics. We evaluate each experiment using multiple metrics used in task-specific benchmarks, such as MSE and MAE for long-horizon forecasting, and sMAPE for short-horizon forecasting. We also note that TimesNet and GPT4TS (Zhou et al., 2023) evaluate anomaly detection performance using vanilla $F_{1}$ score which ignores the sequential nature of time series. Instead, we measure anomaly detection performance with the widely used adjusted best $F_{1}$ score (Goswami et al., 2023a; Challu et al., 2022), and the recently proposed VUS-ROC (Paparrizos et al., 2022a).
Baselines. We compare MOMENT with state-of-the-art deep learning and statistical machine learning models across tasks (Tab. 35). This is in contrast to TimesNet which primarily compared with transformer-based approaches. These comparisons are crucial for assessing the practical utility of the proposed methods. We found that statistical and non-transformer-based approaches like ARIMA for short-horizon forecasting, N-BEATS for long-horizon forecasting, and $k$-nearest neighbors for anomaly detection outperform many deep and transformer-based models.
Hyper-parameter tuning.
[…]
Research questions. Through the following experiments we aim to answer 3 broad research questions.
RQ1: Effectiveness. Is MOMENT effective for multiple time series analysis tasks in limited supervision settings?
RQ2: Interpretability. What is MOMENT learning? Does it capture intuitive time series characteristics such as varying frequencies, trends, and amplitudes?
RQ3: Properties. What is the impact of the size of scaling model size? Can MOMENT, akin to LLMs, be used for cross-modal transfer learning?
$\mathtt{MOMENT_{0}}$ TimesNet GPT4TS TS2Vec T-Loss TNC TS-TCC TST CNN Encoder FCN MCNN MLP ResNet t-LeNet TWIESN DTW
Mean 0.794 0.572 0.566 0.851 0.833 0.786 0.793 0.658 0.751 0.743 0.809 0.702 0.750 0.825 0.348 0.726 0.764
Median 0.815 0.565 0.583 0.871 0.849 0.788 0.802 0.720 0.773 0.753 0.837 0.718 0.766 0.852 0.333 0.724 0.768
Std. 0.147 0.238 0.234 0.134 0.136 0.168 0.176 0.220 0.180 0.159 0.188 0.194 0.169 0.177 0.221 0.164 0.152
Mean rank 7.225 13.324 13.318 3.494 5.261 6.937 6.500 11.846 9.384 8.906 5.565 11.043 9.247 4.406 16.115 10.384 9.071
Median rank 7.0 14.0 14.0 3.0 5.0 6.5 6.0 13.0 9.0 9.0 4.0 12.0 10.0 3.0 17.0 11.0 9.0
Wins/Losses 880.5/566.5 325.5/1121.5 326.0/1121.0 1220.0/227.0 1033.0/375.0 885.5/522.5 924.0/484.0 460.0/987.0 684.0/763.0 727.5/719.5 1031.5/415.5 533.0/914.0 696.5/750.5 1137.0/310.0 71.5/1375.5 593.0/854.0 712.5/734.5
Table 4: Classification accuracy of methods across 91 UCR datasets. Methods with mean and median accuracy higher than MOMENT are in bold. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy. Complete results in Tab. 23.
4.1 MOMENT can solve multiple time series modeling tasks in limited supervision settings
Long-horizon forecasting. Linearly probing MOMENT achieves near state-of-the-art performance on most datasets and horizons, and is only second to PatchTST which generally achieves the lowest MSE (Tab. 2). On many datasets and horizons, forecasting models based on LLMs– TimeLLM and GPT4TS perform worse than MOMENT. Notably, N-BEATS outperforms several recent methods, emphasizing the importance of comparing forecasting performance beyond transformer-based approaches.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(i) Trend (ii) Amplitude (iii) Frequency (iv) Baseline Shift (v) Phase
Figure 4: What is MOMENT learning? Principal components of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and phase information. In each experiment, $c$ controls the factor of interest, for example the power of the trend polynomial $c\in(\frac{1}{8},8)$ (Oreshkin et al., 2020) (Fig. 9), and frequency $c\in(1,32)$ of the generated sine waves (Fig. 9). We generate multiple sine waves by varying $c$, derive their sequence-level representations using MOMENT, and visualize them in a 2- dimensional space using PCA and t-SNE (van der Maaten, 2014) in Fig. 4 and Fig. 7.
Zero-shot short-horizon forecasting. Among all tasks, we found zero-shot short-horizon forecasting to have the largest scope for improvement (Tab. 3). Statistical methods such as Theta and ETS outperformed their deeper counterparts. However, on some datasets, MOMENT achieved lower sMAPE than ARIMA.
Classification. Without any data-specific fine-tuning, MOMENT can learn distinct representations for different classes of data (Fig. 6), and an SVM trained on its representations performs better than all but 4 methods specifically built for time series classification models and trained on each individual dataset. Recently proposed GPT4TS and TimesNet perform poorly despite being trained on each individual dataset with labels.
Model Bit Memory MNIST CIFAR-10 IMDb
GPT-2 1.000 0.975 0.711 0.867
Flan-T5 1.000 0.987 0.672 0.861
MOMENT 1.000 0.982 0.620 0.872
Table 5: Cross-modal transfer experiments. Accuracy measured on the test set, from the checkpoint with the lowest train loss. Even with frozen self-attention and feed-forward layers, MOMENT is able to model cross-modal sequences on par with GPT-2 and Flan-T5 models of similar scale.
Anomaly detection. On 44 time series from the UCR anomaly detection archive, MOMENT consistently outperformed both TimesNet and GPT4TS, as well as 2 state-of-the-art deep learning models tailored for anomaly detection, in both zero-shot and linear probing configurations. However, $k$-nearest neighbors performed marginally better in terms of VUS-ROC score, but had a lower adjusted best $F_{1}$ score.
Dataset $\mathtt{MOMENT_{0}}$ $\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet Naive Linear Nearest Cubic
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 0.082 0.130 0.035 0.075 0.031 0.071 0.036 0.098 0.119 0.108 0.065 0.067 0.083 0.078 0.601 0.153
ETTh1 0.402 0.403 0.139 0.234 0.227 0.254 0.175 0.264 1.185 0.658 0.775 0.534 0.900 0.579 2.178 0.916
ETTh2 0.125 0.238 0.061 0.159 0.109 0.213 0.170 0.286 0.225 0.304 0.135 0.234 0.166 0.252 1.920 0.641
ETTm1 0.202 0.288 0.074 0.168 0.076 0.146 0.087 0.198 0.455 0.365 0.165 0.229 0.230 0.260 0.858 0.494
ETTm2 0.078 0.184 0.031 0.108 0.052 0.133 0.112 0.220 0.113 0.191 0.062 0.138 0.079 0.152 0.534 0.356
Electricity 0.250 0.371 0.094 0.211 0.072 0.183 0.124 0.248 1.474 0.869 0.737 0.592 0.923 0.629 2.257 0.888
Table 6: Imputation Results. MOMENT with linear probing achieved the lowest reconstruction error on all ETT datasets. In the zero-shot setting, MOMENT consistently outperformed all statistical interpolation methods with the exception of linear interpolation. Complete results in Tab. 29.
Metric $\mathtt{MOMENT_{0}}$ $\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet
Anomaly
Transformer
DGHL $k$-NN
Adj. $F_{1}$ Mean 0.585 0.628 0.424 0.537 0.492 0.425 0.554
Median 0.683 0.778 0.331 0.541 0.432 0.331 0.595
Std. 0.377 0.373 0.366 0.389 0.401 0.365 0.393
Mean rank 3.410 3.005 4.862 3.642 4.326 5.071 3.681
Median rank 3.00 3.00 5.00 3.50 4.00 5.25 3.75
Wins/Losses 703.5/472.5 783.0/393.0 419.0/757.0 658.0/518.0 524.0/652.0 378.0/798.0 650.5/525.5
VUS ROC Mean 0.670 0.684 0.611 0.679 0.661 0.646 0.706
Median 0.677 0.692 0.615 0.692 0.658 0.635 0.727
Std. 0.133 0.146 0.114 0.141 0.147 0.137 0.155
Mean rank 4.056 3.382 5.193 3.897 3.913 4.403 3.153
Median rank 4.00 3.00 6.00 4.00 4.00 4.50 3.00
Wins/Losses 577.0/599.0 709.0/467.0 354.0/822.0 608.0/568.0 605.0/571.0 509.0/667.0 754.0/422.0
Table 7: Anomaly detection performance averaged over 248 time series from the UCR Anomaly Archive. $\mathtt{MOMENT_{LP}}$ achieves near state-of-the-art anomaly detection results. Complete results in Tab. 22.
Imputation. Tab. 6 contains imputation performance of all models averaged over 4 different masking rates. MOMENT with linear probing achieved the lowest reconstruction error on all ETT datasets. In the zero-shot setting, MOMENT consistently outperformed all statistical interpolation methods with the exception of linear interpolation.
4.2 What is MOMENT Learning?
We found that MOMENT can capture changes in intuitive time series characteristics such as trend, amplitude, frequencies, and phases of time series. However, it cannot differentiate between vertically shifted time series as it normalizes each signal prior to modeling (Fig. 4,7). Furthermore, on many classification datasets, MOMENT learns distinct representations of different classes, even in a zero-shot setting without access to labels (Fig. 6, 8).
4.3 Properties of Large Time Series Models
Model scaling improves training loss. Like LLMs, we found that increasing the size of the model leads to lower training loss, even before the first epoch (Fig. 6, left). An immediate next step is to assess how effectively this phenomenon extends to time series modeling tasks under limited supervision.
MOMENT can solve cross-modal sequence learning tasks. Lu et al. (2022) first showed that large pre-trained language and vision transformers can solve general sequence learning tasks for modalities outside of text and images with minimal fine-tuning. Several recent studies have leveraged these properties to reprogram LLMs for time series tasks. We explore whether transformers pre-trained on time series can also be used to solve sequence classification tasks on image, text, and binary data. Our results confirm that by freezing the self-attention and feed-forward layers, MOMENT can model sequences comparable to GPT-2 and Flan-T5 models of similar scale (Tab. 5).
MOMENT with randomly initialized weights converges to a lower training loss. Our observations suggest that with sufficient data, pre-training our model from scratch results in a lower training loss than continually pre-training a model of similar size initialized with language modeling weights (Fig. 6, 12). This also underscores that there is sufficient publicly accessible pre-training data available in the Time Series Pile to facilitate pre-training time series foundation models from scratch.
Refer to caption Refer to caption Refer to caption
Figure 5: PCA and t-SNE visualizations of representations learned by MOMENT on the 3 largest UCR datasets. Different colors represent different classes. Even without dataset-specific fine-tuning, MOMENT learns distinct representations for different classes.
Refer to caption Refer to caption
Figure 6: Training losses (MSE). A dashed vertical line denotes the first epoch. All models were trained with a batch size of 131072 patches. (left) Larger models obtain lower training loss. right Eventually, randomly initialized MOMENT-small outperforms the same model initialized with Flan-T5 weights.
5 Conclusion and Future Work
We release the first open-source family of time series foundation models and make contributions at all stages of the development and evaluation process. We first compile a large and diverse collection of public time series, called the Time Series Pile, and demonstrate its efficacy by pre-training high-performing time series foundation models from scratch. Then, we systematically address several time series-specific challenges, which up to now have impeded widespread exploration of extensivelarge-scale multi-dataset pre-training.
We use the Time Series Pile and these strategies to pre-train transformer models of three different sizes. Finally, we design an experimental benchmark to evaluate time series foundation models on multiple practical time series tasks, particularly focusing on scenarios with constrained compute and supervision, building on prior work by Wu et al. (2023). Using this benchmark, we show that MOMENT is effective for the considered tasks with minimal fine-tuning. MOMENT’s superior performance, especially on anomaly detection and classification problems which typically have small datasets, can be attributed to pre-training. Moreover, we demonstrate that across many tasks, smaller statistical and shallower deep learning methods perform reasonably well. Lastly, we make several interesting empirical observations about time series foundation models. Our overarching goal is to push the boundaries of open science by publicly releasing the Time Series Pile, along with code, model weights, and training logs.
We note several interesting directions of future work, including the application of MOMENT to real-world challenges, investigating multi-modal time series and text foundation models (Cai et al., 2023), and enhancing forecasting performance by pre-training MOMENT using causal attention and forecasting objectives.
Acknowledgments
Funding. This work was partially supported by the National Institutes of Health (NIH) under awards R01HL144692 and 1R01NS124642-01, and also by the U.S. Army Research Office and the U.S. Army Futures Command under Contract No. W911NF-20-D-0002. The content of the information does not necessarily reflect the position or the policy of the government and no official endorsement should be inferred.
Discussions. We would like to express our sincerest gratitude to Barış Kurt, Andrey Kan, Laurent Callot, Gauthier Guinet, Jingchao Ni, and Jonas M. Kübler for insightful discussions regarding the problem setting and experimental design. Their unwavering support was instrumental in the development of MOMENT. We are also thankful to Laurent, Barış, Jingchao and Andrey for their constructive feedback on the writing of this manuscript. Additionally, we acknowledge the insightful exchanges with Yuyang (Bernie) Wang, Abdul Fatir Ansari, Ingo Guering, Xiyuan Zhang, and Anoop Deoras. Special thanks to Cherie Ho for suggesting a creative and befitting name for our model. Lastly, we would like to thank Cecilia Morales for her insightful comments, especially on the broader impacts of this work, and for helping us proofread this manuscript.
Data. We extend our gratitude to the authors and data curators whose meticulous efforts were instrumental in curating the datasets utilized for both pre-training and evaluation purposes: UCR Time Series Classification Archive (Dau et al., 2018), TSB-UAD Anomaly Benchmark (Paparrizos et al., 2022b), Monash Forecasting Archive (Godahewa et al., 2021), and the long-horizon forecasting datasets (Zhou et al., 2021).
Software and Models.
Our training and evaluation library was inspired from Time-Series-Library. We would also like to thank the authors of the following libraries for their implementations: universal-computation, Anomaly-Transformer, VUS, tsad-model-selection, One-Fits-All and Statsforecast (Garza et al., 2022).
Reproducibility statement
All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs each with 49 GiB RAM. All MOMENT variants were trained on a single A6000 GPU (with any data or model parallelism). We have made MOMENT-large [^6] https://huggingface.co/AutonLab/MOMENT-1-large and the Time Series Pile [^7] https://huggingface.co/datasets/AutonLab/Timeseries-PILE publicly available on Huggingface. We are working on open-sourcing MOMENT-base and MOMENT-small, and our research code public. The latter is currently available anonymously at https://anonymous.4open.science/r/BETT-773F/README.md. We enlist an exhaustive list of hyper-parameters in App. E to aid reproducibility. We would like to emphasize that all datasets used in this study are publicly available.
Impact statement
Transparency Index. Given the exponential rise in societal reliance on large foundation models, ensuring transparency in their training approach, architecture, and downstream application is crucial for public accountability, scientific advancement, and effective governance. o uphold this objective, we publicly release our training code base, data sources, and evaluation pipeline. We assess the transparency of MOMENT using the criteria outlined by Bommasani et al. (2023), focusing on upstream resources utilized during training and model description, encompassing 32 and 33 transparency indicators, respectively. We report expected upstream and model transparency scores for MOMENT in Tab. 34. Notably, MOMENT is expected to have one of the highest levels of upstream transparency. However, it’s model transparency scores are lower, primarily due to comprehensive (external and third-party) harm and trustworthiness evaluations, which are not well understood in the context of time series modeling.
Environmental Impact. We train multiple models over many days resulting in significant energy usage and a sizeable carbon footprint. However, we hope that releasing our models will ensure that future time series modeling efforts are quicker and more efficient, resulting in lower carbon emissions.
We follow prior work (Bender et al., 2021; Patterson et al., 2021; Touvron et al., 2023; Wu et al., 2022; Dodge et al., 2022) and estimate the carbon footprint of pre-training all variants of MOMENT based on the GPU device used and the carbon efficiency of the electricity grid. Our estimated $\mathtt{CO_{2}}$ generation estimates are shown in Tab. 8.
Model Variant # Parameters (M) GPU Hours Power Consumption (W) Carbon Emission (tCO2eq)
Small 40 308.378 300 31.136
Base 125 308.306 300 31.129
Large 385 404.389 300 40.831
Upper Bound Total - 1021.073 300 103.096
Actual Total - 712.767 300 71.967
Table 8: Total carbon emission induced upon training the MOMENT family of models. MOMENT-small and MOMENT-base were trained simultaneously on a single GPU, thus the TGP required for each model would likely be much less than 300W, and the total time for both models combined is equal to the maximum of the time required for each model. Actual total power consumption and carbon emission values account for this.
We use the Total Graphics Power (TGP) to calculate the total power consumed for training MOMENT models, although the total power consumed by the GPU will likely vary a little based on the GPU utilization while training our model. Our calculations do not account for power demands from other sources of our compute. We use 336.566 Kg $C0_{2}$/MWH as the standard value of $CO_{2}$ emission per megawatt hour of energy consumed for Pittsburgh [^8] https://emissionsindex.org/.
We share an upper limit of the individual $CO_{2}$ emission for each model, as well as a more realistic actual estimate for the carbon emissions from MOMENT-small and MOMENT-base, since they were trained simultaneously on a single Nvidia RTX A6000 GPU, and thus the power consumed by the GPU was shared for the training of both variants. MOMENT-large was trained independently on a single RTX A6000 GPU.
Ethical considerations and potential misuse. Despite MOMENT’s promising performance in limited-data settings, it is important to use its predictions with care, especially in high-stakes settings such as healthcare. Before MOMENT is used for high-stakes decision-making, we recommend fine-tuning and evaluating the model with task-specific in-domain data.
Appendix A Related Work
Transformers and Patching for Time series Modeling. There is a growing body of work utilizing transformers for various time series analysis tasks, for example PatchTST (Nie et al., 2023), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), Pyraformer (Liu et al., 2022) for forecasting; Anomaly Transformer (Xu et al., 2022) for anomaly detection, and TST (Zerveas et al., 2021), TS-TCC (Eldele et al., 2021) for representation learning.
One issue with applying transformers to time series data is the complexity of the self-attention mechanism, which grows quadratically with the size of input tokens (or length of time series). Consequently, the primary focus of most initial applications of transformers to time series, especially for forecasting where longer look-back windows typically improve performance, was to redesign the self-attention mechanism to reduce its complexity (Zhou et al., 2021, 2022; Liu et al., 2022). Nie et al. (2023) demonstrated that treating time series sub-sequences (or patches) as tokens instead of individual time points is a simple, efficient yet effective mechanism for learning useful representations for forecasting. The authors drew inspiration from language and vision domains where sub-words (vs. characters) (Devlin et al., 2019) and 2-D patches (vs. raw pixels) (Bao et al., 2022; Dosovitskiy et al., 2021) are used as inputs to transformers. Drawing inspiration from prior work, we build on top of the transformer architecture which takes disjoint time series sub-sequences (or patches) as input.
Masked Representation Learning. Masked pre-training is a widely-used self-supervised learning task where a model learns to accurately reconstruct masked portions of its input. Masked language (Devlin et al., 2019; Raffel et al., 2020) and image modeling (Xie et al., 2022; Li et al., 2023b) have been successfully utilized to learn models from vast quantities of unlabeled data, which can generalize to a variety of downstream tasks.
For time series data, prior work has primarily focused on contrastive representation learning (Yue et al., 2022; Eldele et al., 2021; Franceschi et al., 2019). The goal of contrastive learning is to learn a representation space where “positive” pairs of time series are close while “negative” pairs are far apart. However, the notion of positive and negative pairs is subjective and data-dependent, and popular transformations such as flipping and cropping invariance may not be appropriate for time series data (Yue et al., 2022). In contrast, some studies mask portions of time series using zeros and learn a model to reconstruct them (Nie et al., 2023; Zerveas et al., 2021; Dong et al., 2023; Li et al., 2023c).
Representation learning via masking is well-suited to all the downstream tasks we care about, especially forecasting and imputation, as they are instances of the masked reconstruction problem. Owing to its simplicity and success in vision and language domains, we use the masked prediction task to pre-train our model, using a special embedding (see [MASK] in Fig. 3) to mask time series patches instead of zeros.
Cross-modal transfer learning using language models. Lu et al. (2022) had first shown that transformers pre-trained on text data (LLMs) can effectively solve sequence modeling tasks in other modalities. Some recent studies have leveraged this inherent ability of language pre-trained transformers to “reprogram” LLMs for time series analysis using parameter efficient fine-tuning and suitable tokenization strategies (Zhou et al., 2023; Gruver et al., 2023; Jin et al., 2023; Cao et al., 2023; Ekambaram et al., 2024). However, some of these models (Jin et al., 2023; Gruver et al., 2023) with billions of parameters demand significant memory and computational resources to perform well. We complement this line of research with three empirical observations (Sec 4.3): we show that (1) transformers trained on time series can also model sequences across modalities, (2) during pre-training, randomly initializing weights lead to lower pre-training loss, than initializing with language modeling weights, and (3) models pre-trained on time series outperform LLM-based models such as (Zhou et al., 2023; Jin et al., 2023) on many tasks and datasets.
Unanswered Questions. To the best of our knowledge, two questions remain largely unanswered in prior work on time series modeling. First, all existing time series models are (pre-)trained and fine-tuned on individual datasets (Nie et al., 2023; Yue et al., 2022; Wu et al., 2023; Zhou et al., 2023), and the benefits (or drawbacks) of large-scale multi-dataset pre-training remains unexplored (Wen et al., 2023). Second, there is very limited work on time series modeling in limited supervision settings, such as zero-shot forecasting (Oreshkin et al., 2021), or few-shot classification (Narwariya et al., 2020). In our work, we consider both these questions and show that pre-training a model of sufficient capacity on a large corpus of unlabeled time series data can in fact enable it to provide reasonably accurate predictions in limited-supervision settings.
Appendix B Contemporary Work
Feature MOMENT Moirai Lag-LLaMa Chronos TimesFM TimeGPT-1 TinyTimeMixers TOTEM
Citation Ours (Woo et al., 2024) (Rasul et al., 2023) (Ansari et al., 2024) (Das et al., 2023) (Garza & Mergenthaler-Canseco, 2023) (Ekambaram et al., 2024) (Talukder et al., 2024)
Base Architecture T5 encoder Encoder-only transformer LLaMa T5 (encoder-decoder) Decoder-only Transformer MLP-Mixer (Tolstikhin et al., 2021) VQ-VAE (Van Den Oord et al., 2017)
Sizes 40M, 125M, 385M 14M, 91M, 311M 200M 20M, 46M, 200M, 710M 17M, 70M, 200M ? $<$1M 0.3M
Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Evaluation Tasks
Forecasting, Classification,
Anomaly detection, Imputation
Forecasting Forecasting Forecasting Forecasting Forecasting Forecasting
Forecasting, Classification,
Anomaly detection, Imputation
Data Total Observations 1.13B++ 27.65B [^9] These values are taken from Tables 8–17 of (Woo et al., 2024) and denote the # of observations 0.36B [^10] These values are taken from Table 4 of (Rasul et al., 2023) and denote the # of observations 84B 100B 100B 244M N/A
# Time Series 13M 4M 11K 11M ? ? N/A N/A
Public $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Public Release Upon acceptance Feb’2024 Feb’2024 Mar’2024 - - Apr’2024 Feb’2024
Tokenization Fixed-length patches Multi-scale patches Lag features Scaling & quantization Fixed-length patches ? Adaptive Patching VQVAE-based
Objective
Reconstruction error (MSE)
Forecast log likelihood
of mixture distribution
Forecast log likelihood
of Student’s t distribution
Forecast Cross-entropy loss Forecasting error (MSE) ? Forecasting error (MSE)
Reconstruction +
Commitment error
Distributional Prediction In progress (post hoc) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ (post hoc) $\times$ $\checkmark$
Interpretability $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
Any-variate Prediction $\checkmark$ (Channel independence) $\checkmark$ (Flattening) $\times$ $\times$ $\times$ $\checkmark$ (?) $\checkmark$ (Channel Independence + Mixing) $\checkmark$ (Flattening)
Context length 512 1000–5000 1024 512 512 ? 512 512
Scaling Model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Data $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Impact of Initialization $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Table 9: Time series foundation modeling has a young but growing literature. Latest ArXiv pre-prints of most contemporary work are from February 2024 with the exception of TimeGPT-1. None of the contemporary works have evaluated their foundation models in terms of cross modal transfer, environmental impact and transparency. ++ denotes that the Time Series Pile is a living database, is still growing in size and diversity. Moirai (Woo et al., 2024) uses a mixture of Student’s t, log-normal, negative binomial, and low variance normal distribution.
Appendix C Interesting directions for future work
We note some interesting directions for future work:
  • Study the impact of design choices such as the impact of the choice of the loss function (Huber, $L_{1}$, $L_{2}$), patch length (4,8)(4,8)( 4 , 8 ), and masking percentage (0.3,0.6)(0.3,0.6)( 0.3 , 0.6 ) on pre-training loss and time series modeling performance.
  • Pre-training data. Two interesting directions include using augmentation and synthetic data (Ansari et al., 2024) to improve the quality of pre-training, and looking at tuning dataset mixtures in the Time Series Pile.
Appendix D The Time Series Pile
We compiled a large collection of publicly available datasets from diverse domains into the Time Series Pile. It has 13 unique domains of data, which includes 20.085 GB worth of 13M unique time series and 1.23 billion timestamps (including channels).
Unique domains Examples of Datasets
Healthcare ECG, EEG, Hospital
Human Body Tongue movement, Finger movement, Muscle Signals
Nature Fish outlines, Flower outlines, River flow
Audio Arabic speech, Japanese Speech, Phonetics
Power Power consumption, Electricity, Home appliance usage
Economics Exchange Rate, Bitcoin, Tourism
Traffic Road Traffic, Pedestrian cross, Line occupancy rate
Weather Temperature, Rain, Wind
Facilities Machine Status, Spacecraft Status, Web traffic
Web Services IOPS, NAB
Synthetic MGAB
Sensors NASA MSL, NASA SMAP
Gait Daphnet
Table 10: The Time Series Pile covers datasets from 13 distinct domains.
Task Dataset Channels Series Length Data Size (Train, Val, Test) Information (Frequency/Number of Classes)
Long
horizon
forecasting
ETTm1, ETTm2 7 {96, 720} (33953, 11425, 11425) Electricity (15 mins)
ETTh1, ETTh2 7 (8033, 2785, 2785) Electricity (15 mins)
Electricity 321 (17805, 2537, 5165) Electricity (Hourly)
Traffic 862 (11673, 1661, 3413) Transportation (Hourly)
Weather 21 (36280, 5175, 10444) Weather (10 mins)
Exchange 8 (4704, 665, 1422) Exchange rate (Daily)
ILI 7 {24, 60} (69, 2, 98) Illness (Weekly)
Short
horizon
forecasting
M4-Yearly 1 6 (16099, 2301, 4600) -
M4-Quarterly 8 (16800, 2400, 4800) -
M4-Monthly 18 (33600, 4800, 9600) -
M3-Yearly 6 (451, 65, 129) -
M3-Quarterly 8 (529, 76, 151) -
M3-Monthly 18 (999, 144, 285) -
Imputation
ETTm1, ETTm2 7 512 (33953, 11425, 11425) Electricity (15 mins)
ETTh1, ETTh2 7 (8033, 2785, 2785) Electricity (15 mins)
Electricity 321 (17805, 2537, 5165) Electricity (Hourly)
Weather 21 (36280, 5175, 10444) Weather (10 mins)
Classification
UWaveGestureLibraryX 1 315 (640, 256, 3582) Motion Gesture (8 classes)
ECG5000 140 (357, 143, 4500) ECG Record (5 classes)
OSULeaf 427 (142, 58, 242) Leaf Outlines (6 classes)
MedicalImages 99 (272, 109, 760) Pixel Intensity (10 classes)
Ham 431 (77, 32, 105) Food spectrographs (2 classes)
Anomaly
detection
(TSB-UAD)
1sddb40 1 - (24489, 9489, 3969) Beats
BIDMC1 - (1274, 204, 7988) PVC
CIMIS44AirTemperature3 - (2346, 632, 3672) Weather Data
CIMIS44AirTemperature5 - (2346, 632, 3672) Weather Data
ECG2 - (10203, 3775, 14488) ECG2 Lead
Table 11: Subset of The Time Series Pile used for experiments. A brief description of datasets that collectively make the Time Series Pile. Due to space constraints, we only include metadata for the subsets of the M3 and M4 datasets in our experiments, as well as 5 classification and anomaly detection datasets. Characteristics of all short-horizon forecasting, classification and anomaly detection datasets in the Time Series Pile can be found in our official repository, and Monash archive, UCR/UEA classification archive, and TSB-UAD anomaly benchmark, respectively.
Dataset Domain # time series Min Length Mean Length Max Length
Total
Observations
Average # of
anomalies
Average # of
abnormal points
Dodger Traffic 1 50399 50399 50399 50399 133 5612
ECG Healthcare 74 5000 311189.82 10828799 23028047 195.6 15634
IOPS Web Services 29 7577 100648.89 149160 2918818 46.5 2312.3
KDD21 Multiple 250 1000 21215.04 250000 5303761 1 196.5
MGAB Synthetic 10 99999 99999 99999 999990 10 200
NAB Web Services 58 1125 6300.72 22693 365442 2 575.5
SensorScope Weather 23 20732 27037.43 30492 621861 11.2 6110.4
YAHOO Web Services 367 740 1560.21 1679 572599 5.9 10.7
NASA-MSL Sensors 27 438 2158.88 4307 58290 1.33 286.3
NASA-SMAP Sensors 54 311 2554.62 2880 137950 1.26 1032.4
Daphnet Gait 45 9599 21759 55039 979155 7.6 2841
GHL Sensors 126 200000 200000 200000 25200000 1.2 388.8
Genesis Sensors 6 16219 16219 16219 97314 3 50
MITDB Healthcare 32 649999 649999 649999 20799968 210.1 72334.3
OPP Sensors 465 22229 31615.85 51115 14701373 2 1267.3
Occupancy Sensors 7 2664 4688.85 9751 32822 18.3 1414.5
SMD Web Services 281 23686 25561.30 28742 7182727 10.4 900.2
SVDB Healthcare 115 230399 230399 230399 26495885 208 27144.5
Table 12: Anomaly detection datasets in Time Series Pile taken from (Paparrizos et al., 2022b). The total number of observations is 129,546,401, while the total number of unique time series is 1,970.
Dataset Domain # Time Series Min Length Mean Length Max Length Total Observations Frequency Forecast Horizon
Australian Electricity Demand Power 5 230736 231052.8 232272 1155264 Half Hourly
Bitcoin Economics 18 2659 4186.88 4581 75364 Daily
Car Parts Economics 2674 51 51 51 136374 monthly
CIF 2016 Economics 72 28 98.72 120 7108 monthly
Covid Deaths Nature 266 212 212 212 56392 Daily
Dominick Economics 115704 28 165.01 393 19092987 weekly
Electricity Power 321 26304 26304 26304 8443584 hourly
321 156 156 156 50076 weekly 8
FRED MD Economics 107 728 728 728 77896 monthly
Hospital Healthcare 767 84 84 84 64428 monthly
Kaggle Web Traffic Web Services 145063 803 803 803 116485589 Daily 59
145063 114 114 114 16537182 weekly 8
KDD Cup 2028 Nature 270 9504 10897.64 10920 2942364 hourly
London Smart Meters Power 5560 288 29951.24 39648 166528896 half_hourly
M1 Multiple 617 48 90.75 150 55998 monthly 18
203 18 48.98 114 9944 quarterly 8
181 15 24.94 58 4515 yearly 6
M3 Multiple 1428 66 117.34 144 167562 monthly 18
174 71 76.58 104 13325 8
756 24 48.94 72 37004 quarterly 8
645 20 28.40 47 18319 yearly 6
M4 Multiple 4227 107 2371.38 9933 10023836 daily 14
414 748 901.86 1008 373372 hourly 48
48000 60 234.30 2812 11246411 monthly 18
24000 24 100.25 874 2406108 quarterly 8
359 93 1035.03 2610 371579 weekly 13
23000 19 37.32 841 858458 yearly 6
NN5 Economics 111 791 791 791 87801 daily 56
111 113 113 113 12543 weekly 8
Pedestrian Counts Traffic 66 576 47459.78 96424 3132346 hourly
Rideshare Traffic 2304 541 541 541 1246464 hourly
Saugeenday Nature 1 23741 23741 23741 23741 daily
Solar Power 137 52560 52560 52560 7200720 10_minutes
1 7397222 7397222 7397222 7397222 4_seconds
137 52 52 52 7124 weekly 5
Sunspot Nature 1 73931 73931 73931 73931 daily
Temperature Rain Nature 32072 725 725 725 23252200 daily
Tourism Economics 366 91 298.57 333 109280 monthly 24
427 30 99.63 130 42544 quarterly 8
518 11 24.62 47 12757 yearly 4
Traffic Traffic 862 17544 17544 17544 15122928 hourly
862 104 104 104 89648 weekly 8
US Births Nature 1 7305 7305 7305 7305 daily
Vehicle Trips Traffic 329 70 128.82 243 42382 daily
Weather Weather 3010 1332 14296.34 65981 43032000 daily
Wind Power 1 7397147 7397147 7397147 7397147 4_seconds
Wind Farms Power 339 6345 507899.88 527040 172178060 minutely
Table 13: Short horizon forecasting datasets in Time Series Pile taken from (Godahewa et al., 2021). The total number of observations is 281,326,601, while the total number of unique time series is 559,102.
Dataset Domain # Time Series Min Length Mean Length Max Length Total Observations Frequency Forecast Horizon
FRED Economics 464 791 1570.08 25289 728520 Daily 14
188028 19 34.76 339 6535989 Yearly 6
1630 94 1261.12 5396 2055634 Weekly 13
133411 60 302.49 3867 40356476 Monthly 18
58908 24 120.11 1288 7075821 Quarterly 8
Table 14: Short horizon forecasting dataset in Time Series Pile taken from the Federal Reserve Economic Data [(FRED)] (Oreshkin et al., 2021). The total number of observations is 56,752,440, while the total number of unique time series is 382,441.
Dataset Domain # Channels Time Series Length Total Observations Frequency Forecast Horizon
Electricity Power 321 26304 8443584 Hourly {96, 192, 336, 720}
ETTh1 Power 7 17420 121940 Hourly {96, 192, 336, 720}
ETTh2 Power 7 17420 121940 Hourly {96, 192, 336, 720}
ETTm1 Power 7 69680 487760 15 Minute {96, 192, 336, 720}
ETTm2 Power 7 69680 487760 15 Minute {96, 192, 336, 720}
Exchange Finance 8 7588 60704 Daily {96, 192, 336, 720}
Illness Epidemiology 7 966 6762 Weekly {24, 36, 48, 60}
Traffic Traffic 862 17544 15122928 Hourly {96, 192, 336, 720}
Weather Weather 21 52696 1106616 10 Minute {96, 192, 336, 720}
Table 15: Long-horizon forecasting datasets in the Time Series Pile taken from Zhou et al. (2021). The total number of observations is 25,959,994, while the total number of unique time series is 1,247.
Table 16: Classification datasets in Time Series Pile taken from (Ismail Fawaz et al., 2019) [UCR Archive]. The total number of observations is 634,084,943, while the total number of unique time series is 290,226.
Dataset Domain
# Time-
series
Classes
Channels
Time Series
Length
# Time Series
(w. channels)
Total
Observations
SemgHandGenderCh2 Human Body 900 2 1 1500 900 1350000
GestureMidAirD2 Human Body 338 26 1 360 338 121680
UWaveGestureLibraryAll Human Body 4478 8 1 945 4478 4231710
SelfRegulationSCP1 Healthcare 561 2 6 896 3366 3015936
UWaveGestureLibraryX Human Body 4478 8 1 315 4478 1410570
GesturePebbleZ2 Human Body 304 6 1 455 304 138320
Healthcare5000 Healthcare 5000 5 1 140 5000 700000
OSULeaf Nature 442 6 1 427 442 188734
MedicalImages Healthcare 1141 10 1 99 1141 112959
Haptics Human Body 463 5 1 1092 463 505596
LargeKitchenAppliances Power 750 3 1 720 750 540000
JapaneseVowels Audio 640 9 12 26 7680 199680
Worms Nature 258 5 1 900 258 232200
Ham Facilities 214 2 1 431 214 92234
DistalPhalanxTW Nature 539 6 1 80 539 43120
ProximalPhalanxOutlineCorrect Nature 891 2 1 80 891 71280
SemgHandMovementCh2 Human Body 900 6 1 1500 900 1350000
RefrigerationDevices Power 750 3 1 720 750 540000
FreezerRegularTrain Facilities 3000 2 1 301 3000 903000
PigAirwayPressure Nature 312 52 1 2000 312 624000
TwoLeadECG Healthcare 1162 2 1 82 1162 95284
GunPointMaleVersusFemale Human Body 451 2 1 150 451 67650
Trace Power 200 4 1 275 200 55000
SmoothSubspace Generated 300 3 1 15 300 4500
MiddlePhalanxTW Nature 553 6 1 80 553 44240
AtrialFibrillation Healthcare 30 3 2 640 60 38400
SyntheticControl Generated 600 6 1 60 600 36000
ShapesAll Generated 1200 60 1 512 1200 614400
Human BodyVerticalSignal Human Body 724 12 1 1250 724 905000
PLAID Facilities 1074 11 1 1344 1074 1443456
AllGestureWiimoteX Human Body 1000 10 1 385 1000 385000
Heartbeat Healthcare 409 2 61 405 24949 10104345
Wafer Facilities 7164 2 1 152 7164 1088928
FaceFour Generated 112 4 1 350 112 39200
Phoneme Audio 2110 39 1 1024 2110 2160640
InlineSkate Human Body 650 7 1 1882 650 1223300
CricketX Human Body 780 12 1 300 780 234000
SelfRegulationSCP2 Healthcare 380 2 7 1152 2660 3064320
DistalPhalanxOutlineCorrect Nature 876 2 1 80 876 70080
ChlorineConcentration Nature 4307 3 1 166 4307 714962
Chinatown Traffic 363 2 1 24 363 8712
GestureMidAirD1 Human Body 338 26 1 360 338 121680
MiddlePhalanxOutlineAgeGroup Nature 554 3 1 80 554 44320
UMD Generated 180 3 1 150 180 27000
Crop Nature 24000 24 1 46 24000 1104000
PenDigits Facilities 10992 10 2 8 21984 175872
GesturePebbleZ1 Human Body 304 6 1 455 304 138320
Handwriting Facilities 1000 26 3 152 3000 456000
Mallat Generated 2400 8 1 1024 2400 2457600
ERing Human Body 300 6 4 65 1200 78000
StarLightCurves Nature 9236 3 1 1024 9236 9457664
WordSynonyms Audio 905 25 1 270 905 244350
PEMS-SF Traffic 440 7 963 144 423720 61015680
FaceDetection Human Body 9414 2 144 62 1355616 84048192
Computers Power 500 2 1 720 500 360000
ArrowHead Generated 211 3 1 251 211 52961
Wine Nature 111 2 1 234 111 25974
Coffee Nature 56 2 1 286 56 16016
Earthquakes Nature 461 2 1 512 461 236032
Herring Nature 128 2 1 512 128 65536
Lightning2 Nature 121 2 1 637 121 77077
Beef Nature 60 5 1 470 60 28200
MiddlePhalanxOutlineCorrect Nature 891 2 1 80 891 71280
HealthcareFiveDays Healthcare 884 2 1 136 884 120224
Yoga Human Body 3300 2 1 426 3300 1405800
Adiac Nature 781 37 1 176 781 137456
HandMovementDirection Human Body 234 4 10 400 2340 936000
MoteStrain Facilities 1272 2 1 84 1272 106848
Rock Nature 70 4 1 2844 70 199080
Strawberry Nature 983 2 1 235 983 231005
InsectWingbeatSound Nature 2200 11 1 256 2200 563200
DodgerLoopWeekend Traffic 158 2 1 288 158 45504
MixedShapesSmallTrain Generated 2525 5 1 1024 2525 2585600
EthanolConcentration Power 524 4 3 1751 1572 2752572
OliveOil Nature 60 4 1 570 60 34200
Meat Nature 120 3 1 448 120 53760
MelbournePedestrian Traffic 3633 10 1 24 3633 87192
Car Facilities 120 4 1 577 120 69240
FaceAll Human Body 2250 14 1 131 2250 294750
FacesUCR Human Body 2250 14 1 131 2250 294750
AllGestureWiimoteY Human Body 1000 10 1 369 1000 369000
NATOPS Human Body 360 6 24 51 8640 440640
SemgHandSubjectCh2 Human Body 900 5 1 1500 900 1350000
ShakeGestureWiimoteZ Human Body 100 10 1 385 100 38500
Cricket Human Body 180 12 6 1197 1080 1292760
BME Generated 180 3 1 128 180 23040
EigenWorms Nature 259 5 6 17984 1554 27947136
FordB Facilities 4446 2 1 500 4446 2223000
NonInvasiveFetalHealthcareThorax1 Healthcare 3765 42 1 750 3765 2823750
UWaveGestureLibrary Human Body 440 8 3 315 1320 415800
CinCHealthcareTorso Healthcare 1420 4 1 1639 1420 2327380
PigArtPressure Nature 312 52 1 2000 312 624000
Fish Nature 350 7 1 463 350 162050
SonyAIBORobotSurface2 Facilities 980 2 1 65 980 63700
FiftyWords Facilities 905 50 1 270 905 244350
MotorImagery Healthcare 378 2 64 3000 24192 72576000
ToeSegmentation1 Human Body 268 2 1 277 268 74236
PhonemeSpectra Audio 6668 39 11 217 73348 15916516
FreezerSmallTrain Facilities 2878 2 1 301 2878 866278
TwoPatterns Generated 5000 4 1 128 5000 640000
ShapeletSim Generated 200 2 1 500 200 100000
Plane Generated 210 7 1 144 210 30240
GestureMidAirD3 Human Body 338 26 1 360 338 121680
DiatomSizeReduction Generated 322 4 1 345 322 111090
Human BodyHorizontalSignal Human Body 724 12 1 1250 724 905000
CricketZ Human Body 780 12 1 300 780 234000
StandWalkJump Human Body 27 3 4 2500 108 270000
WormsTwoClass Human Body 258 2 1 900 258 232200
Lightning7 Nature 143 7 1 319 143 45617
UWaveGestureLibraryY Human Body 4478 8 1 315 4478 1410570
GunPointAgeSpan Human Body 451 2 1 150 451 67650
DistalPhalanxOutlineAgeGroup Nature 539 3 1 80 539 43120
SwedishLeaf Nature 1125 15 1 128 1125 144000
LSST Nature 4925 14 6 36 29550 1063800
CBF Generated 930 3 1 128 930 119040
BeetleFly Nature 40 2 1 512 40 20480
Libras Human Body 360 15 2 45 720 32400
HouseTwenty Facilities 159 2 1 2000 159 318000
ScreenType Facilities 750 3 1 720 750 540000
InsectEPGSmallTrain Nature 266 3 1 601 266 159866
AllGestureWiimoteZ Sensor 1000 10 1 326 1000 326000
DodgerLoopDay Sensor 158 7 1 288 158 45504
NonInvasiveFetalHealthcareThorax2 Healthcare 3765 42 1 750 3765 2823750
BasicHuman Bodys Human Body 80 4 6 100 480 48000
GunPointOldVersusYoung Human Body 451 2 1 150 451 67650
FordA Sensor 4921 2 1 500 4921 2460500
InsectWingbeat Nature 50000 10 200 22 10000000 220000000
ItalyPowerDemand Power 1096 2 1 24 1096 26304
ProximalPhalanxOutlineAgeGroup Nature 605 3 1 80 605 48400
ACSF1 Power 200 10 1 1460 200 292000
GunPoint Human Body 200 2 1 150 200 30000
RacketSports Human Body 303 4 6 30 1818 54540
SmallKitchenAppliances Power 750 3 1 720 750 540000
ProximalPhalanxTW Nature 605 6 1 80 605 48400
DuckDuckGeese Facilities 100 5 1345 270 134500 36315000
PickupGestureWiimoteZ Human Body 100 10 1 361 100 36100
EthanolLevel Power 1004 4 1 1751 1004 1758004
SpokenArabicDigits Audio 8798 10 13 93 114374 10636782
SonyAIBORobotSurface1 Facilities 621 2 1 70 621 43470
HandOutlines Human Body 1370 2 1 2709 1370 3711330
PowerCons Power 360 2 1 144 360 51840
PhalangesOutlinesCorrect Nature 2658 2 1 80 2658 212640
BirdChicken Nature 40 2 1 512 40 20480
ToeSegmentation2 Human Body 166 2 1 343 166 56938
PigCVP Healthcare 312 52 1 2000 312 624000
CricketY Human Body 780 12 1 300 780 234000
FingerMovements Human Body 416 2 28 50 11648 582400
ElectricDevices Power 16637 7 1 96 16637 1597152
InsectEPGRegularTrain Nature 311 3 1 601 311 186911
DodgerLoopGame Traffic 158 2 1 288 158 45504
Fungi Nature 204 18 1 201 204 41004
Symbols Generated 1020 6 1 398 1020 405960
MixedShapesRegularTrain Generated 2925 5 1 1024 2925 2995200
ArticularyWordRecognition Human Body 575 25 9 144 5175 745200
UWaveGestureLibraryZ Human Body 4478 8 1 315 4478 1410570
Epilepsy Human Body 275 4 3 206 825 169950
Healthcare200 Healthcare 200 2 1 96 200 19200
Appendix E Experimental Setup and Results
Through our experiments, our goal is to answer the following research questions.
Is MOMENT effective for multiple time series analysis tasks in limited and rich supervision settings?
We conduct large-scale experiments on widely used benchmarks to evaluate MOMENT on forecasting, classification, anomaly detection, and imputation as outlined in Table 8. The limited supervision setting mimics practical scenarios in which it is infeasible to train (or fine-tune) a deep neural network due to limited compute and, little or inadequately characterized data. In these settings, MOMENT provides predictions without any explicit (re)training on target data [^11] For classification, the quality of MOMENT’s representations is measured using the accuracy of a Support Vector Machine trained on them, as is common in prior work on unsupervised representation learning (Yue et al., 2022; Franceschi et al., 2019). However, unlike prior work, MOMENT embeds time series without any data-specific training.. On the other hand, the rich supervision setting allows us to examine whether MOMENT can utilize task-specific data to improve its performance via end-to-end fine-tuning or linear probing.
What does MOMENT learn?
We evaluated MOMENT’s ability to model time series characteristics such as varying frequencies, trends, and scales. Structure in the PCA and t-SNE (Fig. 9) visualizations of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and auto-correlation information. $\epsilon$ denotes gaussian noise with 0 mean and 0.1 standard deviation. $c$ controls the factor of interest, i.e. the power of the trend polynomial, amplitude, and frequency of the sine waves in experiments (i), (ii) & (iii), respectively.
Hyper-parameter Tuning.
We do not perform extensive hyper-parameter tuning. In all experiments that follow, unless mentioned otherwise, we fine-tune MOMENT-Base with a batch size of 16, and cosine learning rate schedule with an initial learning rate of $1e^{-5}$. For baseline methods, we capture recommended settings from their respective papers and public repositories. We report all hyper-parameters settings for MOMENT and baselines in Appendix D.
E.1 Forecasting
Task description.
Given a time series $\mathcal{T}=[x_{1},...,x_{L}]$ where $x_{i}\in\mathbb{R}$, the univariate forecasting problem is to predict the next $H$ time-steps $[x_{L+1},...,x_{L+H}]$. Depending on the length of the horizon, forecasting can be categorized as short or long-horizon [^12] The distinction between long and short-horizon forecasting is rather arbitrary. For instance, most of the default forecasting horizons for the long-horizon forecasting benchmark Influenza-like Illness (24, 36, 48, 60) are shorter than the Hourly subset of the M4 dataset, a popular short-horizon forecasting benchmark.. We consider both tasks in our experiments. We propose two configurations of MOMENT for the forecasting problem: (1) we can produce short-horizon forecasts without any explicit training or fine-tuning, by appending masked patches and predicting them using the default reconstruction head (Fig. 4 (ii)); (2) alternatively, we can replace the reconstruction head to a forecasting head and then fine-tune it (Fig. 4 (i)).
E.1.1 Long-Horizon Forecasting
Datasets.
We use all the long-horizon forecasting datasets (Sec 3.1). But to speed up our experiments, we drop all exogenous variables from multi-variate datasets and only consider the target time series for forecasting.
Baselines.
We compare our methods with various transformer-based and deep learning baselines. These models can be found in Table 18. For Time-LLM we could not run experiments on Weather, electricity, and traffic datasets, due to time constraints, and since we could not fit them into a single GPU.
Experimental Setting.
We train all models with a look-back window of length $L=$ 512 to forecast $T=$ 24, 60 time-steps for the ILI dataset and $T=$ 96, 720 for the rest. We evaluate the Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics.
Hyperparameters.
The hyperparameters used for training all models in our long-horizon forecasting experiments are shown in Table 17.
Model Hyper-parameters
MOMENT sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.00010.00010.0001
[…]
Time-LLM sequence length : 512512512
patch length : 161616
patch stride length : 888
initial learning rate : 0.0010.0010.001
dimension of feedforward layer : 204820482048
llm layers : 323232
number of heads : 888
N-BEATS sequence length : 512512512
stack types : {trend, seasonality}
number of blocks per stack : 333
[…]
hidden layer units : 256256256
Table 17: Hyper-parameter values for long-horizon forecasting models.
Methods $\mathtt{MOMENT_{LP}}$ Time-LLM GPT4TS PatchTST DLinear TimesNet FEDformer Pyraformer Autoformer Stationary ETSformer LightTS Informer Reformer LogTrans N-BEATS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.154 0.209 - - 0.162 0.212 0.149 0.198 0.176 0.237 0.172 0.220 0.217 0.296 0.896 0.556 0.266 0.336 0.173 0.223 0.197 0.281 0.182 0.242 0.300 0.384 0.689 0.596 0.458 0.490 0.152 0.210
720 0.315 0.336 - - 0.326 0.337 0.314 0.334 0.333 0.362 0.365 0.359 0.403 0.428 1.004 0.934 0.419 0.428 0.414 0.410 0.352 0.288 0.352 0.386 1.059 0.741 1.130 0.792 0.869 0.675 0.331 0.359
ETTh1 96 0.387 0.410 0.408 0.429 0.376 0.397 0.370 0.399 0.375 0.399 0.384 0.402 0.376 0.419 0.664 0.612 0.449 0.459 0.513 0.491 0.494 0.479 0.424 0.432 0.865 0.713 0.837 0.728 0.878 0.740 0.399 0.428
720 0.454 0.472 0.523 0.514 0.477 0.456 0.447 0.466 0.472 0.490 0.521 0.500 0.506 0.507 0.963 0.782 0.514 0.512 0.643 0.616 0.562 0.535 0.547 0.533 1.181 0.865 1.257 0.889 1.135 0.852 0.608 0.573
ETTh2 96 0.288 0.345 0.285 0.348 0.285 0.342 0.274 0.336 0.289 0.353 0.340 0.374 0.358 0.397 0.645 0.597 0.346 0.388 0.476 0.458 0.340 0.391 0.397 0.437 3.755 1.525 2.626 1.317 2.116 1.197 0.327 0.387
720 0.403 0.439 0.399 0.435 0.406 0.441 0.379 0.422 0.605 0.551 0.462 0.468 0.463 0.474 0.963 0.783 0.515 0.511 0.562 0.560 0.500 0.497 0.863 0.672 3.647 1.625 3.874 1.697 3.188 1.540 1.454 0.847
ETTm1 96 0.293 0.349 0.384 0.403 0.292 0.346 0.290 0.342 0.299 0.343 0.338 0.375 0.379 0.419 0.543 0.510 0.505 0.475 0.386 0.398 0.375 0.398 0.374 0.400 0.672 0.571 0.538 0.528 0.600 0.546 0.318 0.367
720 0.405 0.416 0.437 0.429 0.417 0.421 0.416 0.420 0.425 0.421 0.478 0.450 0.543 0.490 0.908 0.724 0.671 0.561 0.585 0.516 0.499 0.462 0.527 0.502 1.166 0.823 1.102 0.841 1.153 0.820 0.448 0.448
ETTm2 96 0.170 0.260 0.181 0.269 0.173 0.262 0.165 0.255 0.167 0.269 0.187 0.267 0.203 0.287 0.435 0.507 0.255 0.339 0.192 0.274 0.189 0.280 0.209 0.308 0.365 0.453 0.658 0.619 0.768 0.642 0.197 0.271
720 0.363 0.387 0.366 0.388 0.378 0.401 0.362 0.385 0.397 0.421 0.408 0.403 0.421 0.415 3.625 1.451 0.433 0.432 0.417 0.413 0.414 0.413 0.675 0.587 3.379 1.338 2.631 1.242 3.048 1.328 0.395 0.419
ILI 24 2.728 1.114 3.025 1.195 2.063 0.881 1.319 0.754 2.215 1.081 2.317 0.934 3.228 1.260 1.420 2.012 3.483 1.287 2.294 0.945 2.527 1.020 8.313 2.144 5.764 1.677 4.400 1.382 4.480 1.444 4.539 1.528
60 2.893 1.132 3.245 1.221 1.979 0.957 1.470 0.788 2.368 1.096 2.027 0.928 2.857 1.157 7.662 2.100 2.770 1.125 2.178 0.963 2.487 1.016 7.283 1.985 5.264 1.564 4.882 1.483 5.278 1.560 5.429 1.661
ECL 96 0.138 0.242 - - 0.139 0.238 0.129 0.222 0.140 0.237 0.168 0.272 0.193 0.308 0.386 0.449 0.201 0.317 0.169 0.273 0.187 0.304 0.207 0.307 0.274 0.368 0.312 0.402 0.258 0.357 0.131 0.228
720 0.211 0.305 - - 0.206 0.297 0.197 0.290 0.203 0.301 0.220 0.320 0.246 0.355 0.376 0.445 0.254 0.361 0.222 0.321 0.233 0.345 0.265 0.360 0.373 0.439 0.340 0.420 0.283 0.376 0.208 0.298
Traffic 96 0.391 0.282 - - 0.388 0.282 0.360 0.249 0.410 0.282 0.593 0.321 0.587 0.366 2.085 0.468 0.613 0.388 0.612 0.338 0.607 0.392 0.615 0.391 0.719 0.391 0.732 0.423 0.684 0.384 0.375 0.259
720 0.450 0.310 - - 0.450 0.312 0.432 0.286 0.466 0.315 0.640 0.350 0.626 0.382 0.881 0.473 0.660 0.408 0.653 0.355 0.632 0.396 0.658 0.407 0.864 0.472 0.755 0.423 0.717 0.396 0.508 0.335
Table 18: Long-term forecasting performance measured using Mean Squared Error (MSE) and Mean Absolute Error (MAE).
Models MOMENT iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear SCINet FEDFormer Stationary Autoformer
Dataset Pred_horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS08 12 0.132 0.249 0.079 0.182 0.133 0.247 0.168 0.232 0.165 0.214 0.227 0.343 0.112 0.212 0.154 0.276 0.087 0.184 0.173 0.273 0.109 0.207 0.436 0.485
24 0.212 0.320 0.115 0.219 0.249 0.343 0.224 0.281 0.215 0.260 0.318 0.409 0.141 0.238 0.248 0.353 0.122 0.221 0.210 0.301 0.140 0.236 0.467 0.502
36 0.309 0.393 0.186 0.235 0.569 0.544 0.321 0.354 0.315 0.355 0.497 0.510 0.198 0.283 0.440 0.470 0.189 0.270 0.320 0.394 0.211 0.294 0.966 0.733
Table 19: Long-term forecasting performance with a look-back window of 96 time-steps. Results are taken from iTransformer (Liu et al., 2023).
E.1.2 Zero-shot Short-Horizon Forecasting
Datasets.
To evaluate zero-shot forecasting performance, we conduct experiments on the M3 and M4 datasets (Sec. 3.1).
Baselines.
We compare MOMENT with GPT4TS (Zhou et al., 2023), TimesNet (Wu et al., 2023), N-BEATS (Oreshkin et al., 2020), 3 statistical and 3 benchmarking forecasting methods: AutoARIMA, AutoTheta, AutoETS, Naive, Seasonal Naive, and Random Walk (Makridakis et al., 2020).
Experimental Setting.
Each statistical method is fit on individual time series before producing a forecast. We follow the same train-test split and forecasting horizons from the M3 and M4 competitions, and report sMAPE as is common in prior work (Oreshkin et al., 2020; Wu et al., 2023) [^13] The definitions of sMAPE were different in the M3 and M4 competitions. In our experiments, we used the same definition as the M4 competition.. We follow the same experimental procedure as outlined in (Oreshkin et al., 2021) with two exceptions: our results are reported only (1) on 40% of the M3 and M4 datasets that were unseen during pre-training, (2) a subset of frequencies with largest support in the datasets. Daily, hourly, and weekly frequencies had very little data and we could not get promising zero-shot performance for any of the deep learning models. Some ways that prior work (Oreshkin et al., 2021) had overcome this issue was by leveraging data from frequencies with plenty of data. We also believe that ensembling played an important part in N-BEATS promising zero-shot performance.
Hyperparameters.
The hyperparameters used for training all models in our short-horizon forecasting experiments are shown in Table 20.
Model Hyper-parameters
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.0020.0020.002
[…]
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial learning rate : 0.0010.0010.001
N-BEATS sequence length : 512512512
stack types : {’trend’, ’seasonality’}
number of blocks per stack : 333
[…]
hidden layer units : 256256256
GPT4TS forecast horizon : 0
gpt layers : 333
patch length : 111
patch stride length : 111
sequence length : 512512512
TimesNet sequence length : 512512512
model dimension : 323232
dimension of feedforward layer : 323232
[…]
Table 20: Hyper-parameter values for short-horizon forecasting models.
Source Dataset $\rightarrow$ M4 Fred
Target Dataset $\downarrow$
M4
Yearly - Yearly
Quarterly - Quarterly
Monthly - Monthly
M3
Yearly Yearly Yearly
Quarterly Quarterly Quarterly
Monthly Monthly Monthly
Table 21: Experimental settings for short-horizon forecasting experiments for varying source and target datasets.
Adjusted Best $F_{1}$ VUS-ROC
Model name Anomaly Transformer $\mathtt{MOMENT_{0}}$ $\mathtt{MOMENT_{LP}}$ DGHL GPT4TS TimesNet AnomalyTransformer $\mathtt{MOMENT_{0}}$ $\mathtt{MOMENT_{LP}}$ DGHL GPT4TS TimesNet
Dataset name
1sddb40 0.030 0.560 0.540 0.390 0.190 0.680 0.640 0.740 0.750 0.640 0.660 0.720
BIDMC1 0.990 1.000 1.000 1.000 1.000 1.000 0.690 0.560 0.650 0.720 0.630 0.740
CHARISfive 0.010 0.070 0.130 0.020 0.020 0.080 0.360 0.430 0.400 0.510 0.450 0.460
CHARISten 0.020 0.060 0.110 0.040 0.100 0.030 0.430 0.500 0.540 0.520 0.510 0.530
CIMIS44AirTemperature3 0.060 1.000 0.980 0.500 0.180 0.470 0.640 0.740 0.750 0.740 0.620 0.740
CIMIS44AirTemperature5 0.390 0.990 0.990 0.960 0.200 0.710 0.780 0.750 0.810 0.920 0.560 0.720
ECG2 1.000 1.000 1.000 0.620 0.900 1.000 0.830 0.740 0.840 0.630 0.780 0.600
ECG3 0.360 0.810 0.980 0.800 0.840 0.480 0.540 0.700 0.770 0.680 0.450 0.610
Fantasia 0.750 1.000 0.950 0.660 0.870 0.550 0.730 0.630 0.640 0.710 0.650 0.610
GP711MarkerLFM5z4 0.930 0.810 1.000 0.500 0.640 0.950 0.540 0.630 0.730 0.600 0.620 0.720
GP711MarkerLFM5z5 0.760 0.690 0.970 0.310 0.480 0.900 0.690 0.760 0.720 0.520 0.630 0.840
InternalBleeding4 NaN 1.000 NaN NaN NaN NaN NaN 0.650 NaN NaN NaN NaN
InternalBleeding5 0.940 1.000 1.000 1.000 0.920 1.000 0.460 0.600 0.690 0.760 0.630 0.940
Italianpowerdemand 0.010 0.390 0.740 0.590 0.010 0.440 0.450 0.800 0.770 0.700 0.480 0.710
Lab2Cmac011215EPG5 0.990 0.970 0.980 0.340 0.600 0.990 0.770 0.620 0.630 0.710 0.640 0.610
Lab2Cmac011215EPG6 0.410 0.090 0.100 0.260 0.100 0.170 0.700 0.480 0.480 0.600 0.520 0.450
MesoplodonDensirostris 1.000 0.910 0.840 0.790 1.000 1.000 0.850 0.730 0.720 0.740 0.690 0.790
PowerDemand1 0.870 0.260 0.440 0.490 0.760 0.950 0.720 0.520 0.540 0.530 0.600 0.750
TkeepFirstMARS 0.010 0.080 0.150 0.020 0.020 0.230 0.520 0.570 0.760 0.460 0.500 0.790
TkeepSecondMARS 0.830 0.950 1.000 0.160 0.120 0.950 0.720 0.950 0.910 0.970 0.810 0.980
WalkingAceleration5 0.990 1.000 1.000 0.910 0.870 0.930 0.940 0.860 0.870 0.930 0.910 0.850
apneaecg 0.400 0.210 0.200 0.250 0.310 0.260 0.580 0.690 0.690 0.590 0.580 0.760
apneaecg2 0.650 0.940 1.000 1.000 1.000 0.650 0.790 0.750 0.740 0.730 0.650 0.610
gait1 0.180 0.710 0.360 0.070 0.410 0.520 0.630 0.650 0.570 0.600 0.580 0.600
gaitHunt1 0.080 0.500 0.430 0.020 0.100 0.300 0.810 0.640 0.680 0.570 0.710 0.840
insectEPG2 0.120 0.110 0.230 0.140 0.810 0.960 0.650 0.570 0.820 0.650 0.560 0.730
insectEPG4 0.980 1.000 1.000 0.460 0.210 0.850 0.690 0.700 0.720 0.730 0.490 0.650
ltstdbs30791AS 1.000 1.000 1.000 1.000 1.000 1.000 0.780 0.760 0.810 0.770 0.740 0.670
mit14046longtermecg 0.450 0.560 0.590 0.530 0.580 0.600 0.790 0.660 0.660 0.640 0.610 0.840
park3m 0.150 0.560 0.640 0.200 0.630 0.930 0.630 0.750 0.780 0.650 0.540 0.780
qtdbSel1005V 0.410 0.570 0.650 0.400 0.390 0.530 0.520 0.640 0.640 0.490 0.610 0.540
qtdbSel100MLII 0.420 0.780 0.840 0.410 0.600 0.870 0.620 0.580 0.620 0.590 0.580 0.650
resperation1 0.000 0.040 0.150 0.030 0.010 0.030 0.750 0.500 0.670 0.740 0.470 0.670
s20101mML2 0.690 0.650 0.710 0.150 0.050 0.080 0.640 0.760 0.720 0.690 0.640 0.690
sddb49 0.890 1.000 1.000 0.880 0.940 1.000 0.660 0.730 0.730 0.740 0.580 0.680
sel840mECG1 0.160 0.610 0.660 0.280 0.210 0.360 0.620 0.720 0.720 0.870 0.650 0.600
sel840mECG2 0.150 0.360 0.390 0.320 0.280 0.210 0.590 0.710 0.690 0.490 0.520 0.520
tilt12744mtable 0.070 0.110 0.240 0.100 0.000 0.030 0.480 0.670 0.740 0.660 0.510 0.640
tilt12754table 0.230 0.590 0.640 0.040 0.060 0.050 0.600 0.750 0.820 0.790 0.550 0.750
tiltAPB2 0.920 0.960 0.980 0.360 0.830 0.380 0.770 0.750 0.770 0.710 0.600 0.700
tiltAPB3 0.170 0.480 0.850 0.030 0.050 0.090 0.680 0.610 0.650 0.540 0.440 0.580
weallwalk 0.000 0.520 0.580 0.070 0.130 0.170 0.730 0.930 0.930 0.860 0.870 0.850
Table 22: Anomaly detection performance measured using adj. best $F_{1}$ and VUS-ROC for a subset of 45 datasets sampled from the UCR Anomaly archive.
Dataset $\mathtt{MOMENT_{0}}$ TimesNet GPT4TS TS2Vec T-Loss TNC TS-TCC TST CNN Encoder FCN MCDNN MLP ResNet t-LeNet TWIESN DTW
GestureMidAirD2 0.608 0.131 0.200 0.469 0.546 0.254 0.254 0.138 0.518 0.480 0.631 0.500 0.545 0.668 0.038 0.575 0.608
UWaveGestureLibraryX 0.821 0.688 0.749 0.795 0.785 0.733 0.733 0.569 0.721 0.771 0.754 0.726 0.768 0.781 0.127 0.608 0.728
GesturePebbleZ2 0.816 0.310 0.285 0.873 0.899 0.430 0.430 0.380 0.778 0.796 0.781 0.720 0.701 0.777 0.184 0.843 0.671
ECG5000 0.942 0.584 0.584 0.935 0.933 0.941 0.941 0.928 0.928 0.941 0.940 0.933 0.930 0.935 0.584 0.922 0.924
OSULeaf 0.785 0.397 0.231 0.851 0.760 0.723 0.723 0.545 0.482 0.554 0.979 0.419 0.560 0.980 0.182 0.628 0.591
MedicalImages 0.762 0.571 0.496 0.789 0.750 0.747 0.747 0.632 0.671 0.664 0.778 0.627 0.719 0.770 0.514 0.649 0.737
Ham 0.581 0.686 0.781 0.714 0.724 0.743 0.743 0.524 0.720 0.682 0.707 0.718 0.699 0.758 0.514 0.768 0.467
DistalPhalanxTW 0.612 0.604 0.619 0.698 0.676 0.676 0.676 0.568 0.671 0.694 0.695 0.685 0.610 0.663 0.285 0.591 0.590
ProximalPhalanxOutlineCorrect 0.856 0.869 0.801 0.887 0.859 0.873 0.873 0.770 0.807 0.768 0.907 0.866 0.730 0.920 0.684 0.817 0.784
FreezerRegularTrain 0.982 0.926 0.829 0.986 0.956 0.989 0.989 0.922 0.987 0.760 0.997 0.973 0.906 0.998 0.500 0.946 0.899
TwoLeadECG 0.847 0.633 0.658 0.986 0.999 0.976 0.976 0.871 0.877 0.784 0.999 0.806 0.753 1.000 0.500 0.949 0.905
GunPointMaleVersusFemale 0.991 0.601 0.475 1.000 0.997 0.997 0.997 1.000 0.977 0.978 0.997 0.952 0.980 0.992 0.525 0.988 0.997
Trace 1.000 0.760 0.710 1.000 0.990 1.000 1.000 1.000 0.952 0.740 1.000 0.902 0.806 1.000 0.240 0.934 1.000
SmoothSubspace 0.820 0.440 0.453 0.980 0.960 0.953 0.953 0.827 0.976 0.964 0.975 0.963 0.980 0.980 0.333 0.849 0.827
MiddlePhalanxTW 0.532 0.506 0.571 0.584 0.591 0.610 0.610 0.506 0.551 0.597 0.501 0.562 0.536 0.495 0.286 0.569 0.506
SyntheticControl 0.990 0.467 0.437 0.997 0.987 0.990 0.990 0.490 0.987 0.973 0.989 0.953 0.973 0.997 0.167 0.879 0.993
ShapesAll 0.815 0.238 0.237 0.902 0.848 0.773 0.773 0.733 0.617 0.679 0.894 0.599 0.776 0.926 0.017 0.643 0.768
AllGestureWiimoteX 0.607 0.209 0.237 0.777 0.763 0.697 0.697 0.259 0.411 0.475 0.713 0.261 0.477 0.741 0.100 0.522 0.716
Wafer 0.997 0.989 0.994 0.998 0.992 0.994 0.994 0.991 0.961 0.998 0.997 0.992 0.996 0.998 0.892 0.916 0.980
FaceFour 0.852 0.830 0.659 0.932 0.920 0.773 0.773 0.511 0.905 0.852 0.930 0.711 0.836 0.955 0.295 0.857 0.830
CricketX 0.749 0.523 0.531 0.782 0.713 0.731 0.731 0.385 0.535 0.644 0.794 0.513 0.591 0.799 0.074 0.627 0.754
DistalPhalanxOutlineCorrect 0.717 0.786 0.659 0.761 0.775 0.754 0.754 0.728 0.772 0.724 0.760 0.759 0.727 0.770 0.583 0.711 0.717
ChlorineConcentration 0.765 0.618 0.565 0.832 0.749 0.753 0.753 0.562 0.608 0.583 0.817 0.662 0.800 0.853 0.533 0.554 0.648
Chinatown 0.965 0.274 0.857 0.965 0.951 0.983 0.983 0.936 0.977 0.966 0.980 0.945 0.872 0.978 0.726 0.825 0.957
GestureMidAirD1 0.646 0.285 0.292 0.608 0.608 0.369 0.369 0.208 0.534 0.528 0.695 0.518 0.575 0.698 0.038 0.549 0.569
MiddlePhalanxOutlineAgeGroup 0.461 0.344 0.526 0.636 0.656 0.630 0.630 0.617 0.534 0.577 0.535 0.558 0.522 0.545 0.571 0.578 0.500
UMD 0.993 0.681 0.368 1.000 0.993 0.986 0.986 0.910 0.960 0.771 0.988 0.842 0.949 0.990 0.333 0.835 0.993
Crop 0.734 0.388 0.341 0.756 0.722 0.742 0.742 0.710 0.670 0.760 0.738 0.687 0.618 0.743 0.042 0.489 0.665
GesturePebbleZ1 0.849 0.512 0.605 0.930 0.919 0.395 0.395 0.500 0.844 0.821 0.880 0.769 0.792 0.901 0.163 0.840 0.791
WordSynonyms 0.688 0.335 0.451 0.676 0.691 0.531 0.531 0.422 0.568 0.557 0.561 0.470 0.599 0.617 0.219 0.506 0.649
ArrowHead 0.743 0.360 0.429 0.857 0.766 0.737 0.737 0.771 0.717 0.630 0.843 0.678 0.784 0.838 0.303 0.689 0.703
Wine 0.537 0.519 0.611 0.870 0.815 0.778 0.778 0.500 0.519 0.556 0.611 0.500 0.541 0.722 0.500 0.744 0.574
Coffee 0.893 0.964 0.679 1.000 1.000 1.000 1.000 0.821 1.000 0.886 1.000 0.979 0.993 1.000 0.507 0.979 1.000
Earthquakes 0.748 0.741 0.748 0.748 0.748 0.748 0.748 0.748 0.709 0.740 0.725 0.748 0.727 0.712 0.748 0.748 0.719
Herring 0.594 0.531 0.578 0.641 0.594 0.594 0.594 0.594 0.531 0.512 0.644 0.572 0.491 0.600 0.594 0.625 0.531
Beef 0.833 0.400 0.167 0.767 0.667 0.600 0.600 0.500 0.767 0.707 0.680 0.507 0.713 0.753 0.200 0.527 0.633
MiddlePhalanxOutlineCorrect 0.467 0.512 0.519 0.838 0.825 0.818 0.818 0.753 0.744 0.752 0.795 0.796 0.755 0.826 0.570 0.743 0.698
ECGFiveDays 0.804 0.519 0.561 1.000 1.000 0.878 0.878 0.763 0.874 0.842 0.985 0.800 0.973 0.966 0.497 0.723 0.768
Yoga 0.834 0.672 0.691 0.887 0.837 0.791 0.791 0.830 0.786 0.753 0.837 0.741 0.856 0.867 0.536 0.626 0.837
Adiac 0.688 0.565 0.598 0.762 0.675 0.767 0.767 0.550 0.393 0.318 0.841 0.620 0.391 0.833 0.023 0.428 0.604
MoteStrain 0.774 0.700 0.681 0.861 0.851 0.843 0.843 0.768 0.885 0.872 0.936 0.691 0.855 0.924 0.539 0.809 0.835
Strawberry 0.951 0.946 0.935 0.962 0.954 0.965 0.965 0.916 0.952 0.959 0.975 0.958 0.959 0.980 0.643 0.911 0.941
InsectWingbeatSound 0.607 0.529 0.598 0.630 0.597 0.415 0.415 0.266 0.585 0.630 0.392 0.587 0.604 0.499 0.091 0.435 0.355
DodgerLoopWeekend 0.826 0.638 0.804 0.964 NaN NaN NaN 0.732 0.974 0.983 0.904 0.978 0.978 0.952 0.739 0.954 0.949
Meat 0.917 0.433 0.667 0.950 0.950 0.883 0.883 0.900 0.913 0.787 0.803 0.787 0.893 0.990 0.333 0.970 0.933
MelbournePedestrian 0.876 0.718 0.207 0.959 0.944 0.949 0.949 0.741 0.813 0.884 0.912 0.840 0.863 0.909 0.100 0.730 0.791
FaceAll 0.791 0.177 0.147 0.771 0.786 0.813 0.813 0.504 0.774 0.794 0.938 0.720 0.794 0.867 0.080 0.673 0.808
FacesUCR 0.811 0.679 0.462 0.924 0.884 0.863 0.863 0.543 0.873 0.867 0.943 0.775 0.831 0.954 0.143 0.641 0.905
AllGestureWiimoteY 0.666 0.223 0.160 0.793 0.726 0.741 0.741 0.423 0.479 0.509 0.784 0.420 0.571 0.794 0.100 0.600 0.729
ShakeGestureWiimoteZ 0.960 0.020 0.080 0.940 0.920 0.860 0.860 0.760 0.580 0.756 0.884 0.516 0.548 0.880 0.100 0.864 0.860
BME 0.960 0.467 0.367 0.993 0.993 0.933 0.933 0.760 0.947 0.827 0.836 0.896 0.905 0.999 0.333 0.819 0.900
FordB 0.798 0.754 0.677 0.794 0.793 0.815 0.815 0.507 0.749 0.777 0.772 0.698 0.707 0.813 0.503 0.512 0.620
Fish 0.800 0.726 0.731 0.926 0.891 0.817 0.817 0.720 0.855 0.734 0.961 0.720 0.848 0.981 0.126 0.878 0.823
SonyAIBORobotSurface2 0.829 0.646 0.650 0.871 0.889 0.907 0.907 0.745 0.831 0.844 0.980 0.804 0.831 0.975 0.617 0.635 0.831
FiftyWords 0.802 0.499 0.492 0.771 0.732 0.653 0.653 0.525 0.624 0.658 0.646 0.611 0.708 0.740 0.125 0.518 0.690
ToeSegmentation1 0.925 0.456 0.561 0.917 0.939 0.930 0.930 0.807 0.598 0.706 0.961 0.559 0.589 0.957 0.526 0.882 0.772
FreezerSmallTrain 0.902 0.704 0.500 0.870 0.933 0.979 0.979 0.920 0.739 0.676 0.683 0.688 0.686 0.832 0.500 0.917 0.753
TwoPatterns 0.994 0.989 0.923 1.000 0.999 0.999 0.999 0.466 0.991 1.000 0.870 0.976 0.948 1.000 0.259 0.875 1.000
ShapeletSim 0.961 0.500 0.489 1.000 0.672 0.683 0.683 0.489 0.497 0.510 0.706 0.498 0.513 0.782 0.500 0.546 0.650
Plane 0.990 0.981 0.924 1.000 0.990 1.000 1.000 0.933 0.962 0.964 1.000 0.952 0.977 1.000 0.143 1.000 1.000
GestureMidAirD3 0.369 0.085 0.162 0.292 0.285 0.177 0.177 0.154 0.317 0.368 0.326 0.278 0.382 0.340 0.038 0.275 0.323
DiatomSizeReduction 0.879 0.967 0.987 0.984 0.984 0.977 0.977 0.961 0.954 0.880 0.346 0.646 0.909 0.301 0.301 0.914 0.967
CricketZ 0.731 0.459 0.397 0.792 0.708 0.713 0.713 0.403 0.501 0.651 0.810 0.484 0.629 0.809 0.062 0.643 0.754
Lightning7 0.726 0.575 0.562 0.863 0.795 0.685 0.685 0.411 0.647 0.696 0.825 0.559 0.616 0.827 0.260 0.608 0.726
UWaveGestureLibraryY 0.738 0.547 0.648 0.719 0.710 0.641 0.641 0.348 0.626 0.676 0.642 0.639 0.699 0.666 0.121 0.497 0.634
GunPointAgeSpan 0.962 0.494 0.494 0.987 0.994 0.994 0.994 0.991 0.912 0.890 0.996 0.887 0.934 0.997 0.494 0.965 0.918
DistalPhalanxOutlineAgeGroup 0.669 0.597 0.489 0.727 0.727 0.755 0.755 0.741 0.758 0.761 0.718 0.729 0.647 0.718 0.433 0.705 0.770
SwedishLeaf 0.923 0.894 0.899 0.941 0.914 0.923 0.923 0.738 0.884 0.902 0.967 0.841 0.845 0.963 0.064 0.837 0.792
CBF 0.960 0.761 0.830 1.000 0.983 0.998 0.998 0.898 0.959 0.977 0.994 0.908 0.869 0.996 0.332 0.896 0.997
BeetleFly 0.900 0.400 0.700 0.900 0.800 0.800 0.800 1.000 0.900 0.620 0.910 0.630 0.880 0.850 0.500 0.790 0.700
AllGestureWiimoteZ 0.537 0.221 0.116 0.746 0.723 0.689 0.689 0.447 0.375 0.396 0.692 0.287 0.439 0.726 0.100 0.516 0.643
DodgerLoopDay 0.438 0.237 0.200 0.562 NaN NaN NaN 0.200 0.312 0.487 0.143 0.305 0.160 0.150 0.160 0.593 0.500
GunPointOldVersusYoung 0.981 0.508 0.524 1.000 1.000 1.000 1.000 1.000 0.922 0.923 0.989 0.926 0.941 0.989 0.524 0.975 0.838
FordA 0.936 0.913 0.914 0.936 0.928 0.930 0.930 0.568 0.896 0.928 0.914 0.863 0.816 0.937 0.510 0.555 0.555
ItalyPowerDemand 0.911 0.837 0.880 0.925 0.954 0.955 0.955 0.845 0.954 0.964 0.963 0.966 0.953 0.962 0.499 0.871 0.950
ProximalPhalanxOutlineAgeGroup 0.863 0.868 0.839 0.834 0.844 0.839 0.839 0.854 0.812 0.872 0.825 0.839 0.849 0.847 0.488 0.839 0.805
GunPoint 0.927 0.887 0.847 0.980 0.980 0.993 0.993 0.827 0.948 0.784 1.000 0.907 0.928 0.991 0.493 0.989 0.907
ProximalPhalanxTW 0.712 0.800 0.712 0.824 0.771 0.800 0.800 0.780 0.777 0.791 0.761 0.775 0.767 0.773 0.341 0.784 0.761
PickupGestureWiimoteZ 0.620 0.100 0.080 0.820 0.740 0.600 0.600 0.240 0.608 0.496 0.744 0.412 0.604 0.704 0.100 0.616 0.660
SonyAIBORobotSurface1 0.729 0.542 0.589 0.903 0.902 0.899 0.899 0.724 0.690 0.729 0.958 0.655 0.692 0.961 0.429 0.725 0.725
PowerCons 0.894 0.956 0.989 0.961 0.900 0.961 0.961 0.911 0.960 0.971 0.863 0.929 0.977 0.879 0.500 0.852 0.878
PhalangesOutlinesCorrect 0.652 0.614 0.663 0.809 0.784 0.804 0.804 0.773 0.799 0.745 0.818 0.795 0.756 0.845 0.613 0.656 0.728
BirdChicken 0.850 0.450 0.550 0.800 0.850 0.650 0.650 0.650 0.710 0.510 0.940 0.540 0.740 0.880 0.500 0.620 0.750
ToeSegmentation2 0.915 0.731 0.731 0.892 0.900 0.877 0.877 0.615 0.752 0.702 0.889 0.649 0.745 0.894 0.815 0.794 0.838
CricketY 0.746 0.531 0.521 0.749 0.728 0.718 0.718 0.467 0.582 0.639 0.793 0.521 0.598 0.810 0.085 0.652 0.744
ElectricDevices 0.646 0.552 0.506 0.721 0.707 0.686 0.686 0.676 0.686 0.702 0.706 0.653 0.593 0.728 0.242 0.605 0.602
DodgerLoopGame 0.623 0.471 0.717 0.841 NaN NaN NaN 0.696 0.816 0.810 0.768 0.877 0.865 0.710 0.478 0.716 0.877
Fungi 0.898 0.043 0.054 0.957 1.000 0.753 0.753 0.366 0.961 0.934 0.018 0.051 0.863 0.177 0.063 0.439 0.839
Symbols 0.936 0.864 0.694 0.976 0.963 0.916 0.916 0.786 0.808 0.754 0.955 0.644 0.836 0.893 0.174 0.798 0.950
UWaveGestureLibraryZ 0.765 0.632 0.643 0.770 0.757 0.690 0.690 0.655 0.630 0.684 0.727 0.645 0.697 0.749 0.121 0.573 0.658
ECG200 0.760 0.830 0.790 0.920 0.940 0.880 0.880 0.830 0.816 0.884 0.888 0.838 0.914 0.874 0.640 0.874 0.770
Table 23: Classification accuracy of methods across 91 UCR datasets. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy.
Dataset MOMENT0 TS2Vec T-Loss TNC TS-TCC TST DTW
ArticularyWordRecognition 0.990 0.987 0.943 0.973 0.953 0.977 0.987
AtrialFibrillation 0.200 0.200 0.133 0.133 0.267 0.067 0.200
BasicMotions 1.000 0.975 1.000 0.975 1.000 0.975 0.975
Cricket 0.986 0.972 0.972 0.958 0.917 1.000 1.000
DuckDuckGeese 0.600 0.680 0.650 0.460 0.380 0.620 0.600
EigenWorms 0.809 0.847 0.840 0.840 0.779 0.748 0.618
Epilepsy 0.993 0.964 0.971 0.957 0.957 0.949 0.964
ERing 0.959 0.874 0.133 0.852 0.904 0.874 0.133
EthanolConcentration 0.357 0.308 0.205 0.297 0.285 0.262 0.323
FaceDetection 0.633 0.501 0.513 0.536 0.544 0.534 0.529
FingerMovements 0.490 0.480 0.580 0.470 0.460 0.560 0.530
HandMovementDirection 0.324 0.338 0.351 0.324 0.243 0.243 0.231
Handwriting 0.308 0.515 0.451 0.249 0.498 0.225 0.286
Heartbeat 0.722 0.683 0.741 0.746 0.751 0.746 0.717
JapaneseVowels 0.716 0.984 0.989 0.978 0.930 0.978 0.949
Libras 0.850 0.867 0.883 0.817 0.822 0.656 0.870
LSST 0.411 0.537 0.509 0.595 0.474 0.408 0.551
MotorImagery 0.500 0.510 0.580 0.500 0.610 0.500 0.500
NATOPS 0.828 0.928 0.917 0.911 0.822 0.850 0.883
PEMS-SF 0.896 0.682 0.676 0.699 0.734 0.740 0.711
PenDigits 0.972 0.989 0.981 0.979 0.974 0.560 0.977
PhonemeSpectra 0.233 0.233 0.222 0.207 0.252 0.085 0.151
RacketSports 0.796 0.855 0.855 0.776 0.816 0.809 0.803
SelfRegulationSCP1 0.840 0.812 0.843 0.799 0.823 0.754 0.775
SelfRegulationSCP2 0.478 0.578 0.539 0.550 0.533 0.550 0.539
SpokenArabicDigits 0.981 0.988 0.905 0.934 0.970 0.923 0.963
StandWalkJump 0.400 0.467 0.333 0.400 0.333 0.267 0.200
UWaveGestureLibrary 0.909 0.906 0.875 0.759 0.753 0.575 0.903
InsectWingbeat 0.246 0.466 0.156 0.469 0.264 0.105 NaN
Mean 0.670 0.694 0.646 0.660 0.657 0.605 0.638
Median 0.722 0.683 0.676 0.746 0.751 0.620 0.664
Std. 0.274 0.255 0.296 0.267 0.263 0.294 0.296
Mean Rank 3.466 2.862 3.603 4.362 4.121 5.069 4.429
Median Rank 3.0 2.5 4.0 5.0 5.0 5.5 4.5
Wins/Losses 101.5/71.5 119.0/54.0 97.5/75.5 75.5/97.5 82.5/90.5 55.0/118.0 72.0/96.0
Table 24: Classification accuracy of methods across 29 UEA datasets. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy.
E.2 Classification
Task Description.
The classification problem comprises of learning a mapping $f:\mathcal{T}\to\{1,...,C\}$ from a time series to a finite set of classes, using a training dataset of the form $\{(\mathcal{T}_{0},c_{0}),...,(\mathcal{T}_{n},c_{n})\}$, $c_{i}\in\{1,...,C\}$. One straightforward way to use MOMENT to learn $f$ is to replace its reconstruction head with a linear head that maps patch representations to the $C$ logits. Another way would be to learn $f$ in two stages, as is common in prior work on unsupervised representation learning (Yue et al., 2022; Franceschi et al., 2019): in the first stage, we obtain sequence-level representations for each time series without access to labels. The second stage involves learning any ML classifier (e.g., Support Vector Machine with RBF kernel) using these representations and labels.
Datasets.
We conduct experiments on a subset of 95 datasets from the UCR Classification Archive (Dau et al., 2018). These datasets (listed in Table 10) comprise of equal-length univariate time series shorter than 512 time steps.
Baselines.
We compare MOMENT against 5 unsupervised representation learning methods (TS2Vec (Yue et al., 2022), TST (Zerveas et al., 2021), TS-TCC (Eldele et al., 2021), TNC (Tonekaboni et al., 2021), and T-Loss (Franceschi et al., 2019)), 8 supervised deep learning (CNN (Zebik et al., 2017), Encoder (Serrà et al., 2018), FCN (Wang et al., 2017), MCNN (Cui et al., 2016), MLP (Wang et al., 2017), ResNet (Wang et al., 2017), t-LeNet (Le Guennec et al., 2016), TWIESN (Tanisaro & Heidemann, 2016)), 1 supervised statistical learning method DTW (Dau et al., 2018)), TimesNet (Wu et al., 2023) and GPT4TS (Zhou et al., 2023).
Experimental Setting.
All models except for MOMENT were trained on each dataset individually, either with labels for supervised deep and statistical learning methods), or without labels for representation learning methods. We collect baseline results for deep learning methods from Ismail Fawaz et al. (2019), representation learning methods from Yue et al. (2022), and DTW from Dau et al. (2018). We report accuracy as the evaluation metric.
Hyperparameters.
The hyperparameters used for evaluating classification experiments are shown in Table 25.
Model Hyper-parameters
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
[…]
kernel : RBF
degree : 333
cache size : 200200200
max iterations : 100000001000000010000000
decision function shape : One versus rest
Table 25: Hyper-parameter values for classification.
E.3 Anomaly Detection
Task Description.
Given a time series $\mathcal{T}$, anomaly detection is a binary classification problem, where the goal is to detect whether a time step $x_{i}$ is indicative of an anomaly or not. As shown in Fig. 4 (v), to detect anomalies in $\mathcal{T}$, we retain MOMENT’s reconstruction head and use it to reconstruct the input time series. Then, time steps where observations and predictions differ beyond a certain threshold are classified as anomalies [^14] Estimating good thresholds for anomaly detection is beyond the scope of this study and an active area of research (Goswami et al., 2023a; Schmidl et al., 2022)..
Datasets.
We conduct experiments on a subset of 46 univariate time series from the UCR Anomaly Archive (Wu & Keogh, 2023), as enumerated in Table 11. When choosing the subset of time series, we prioritized coverage over different domains and data sources represented in the archive.
Baselines.
We compare MOMENT with 2 state-of-the-art anomaly detection methods DGHL (Challu et al., 2022) and Anomaly Transformer (Xu et al., 2022) along with TimesNet and GPT4TS. We also include $k$-Nearest Neighbors (with $k=$ 5) (Ramaswamy et al., 2000), a classical anomaly detection method in our experiments. In the zero-shot setting, we compare MOMENT to randomly initialized DGHL ($\mathtt{DGHL_{0}}$) [^15] Randomly initialized DGHL is not a trivial zero-shot baseline, since it performs gradient descent to find the best latent $z$ that minimizes reconstruction error during inference time (Challu et al., 2022). and $k$-NN.
Experimental Setting.
All algorithms use a fixed anomaly detection window size (= 512). Based on prior work (Wu et al., 2023; Zhou et al., 2023), we use the mean squared error between predictions and observations as the anomaly criterion [^16] To ensure a fair comparison, we do not use Anomaly Transformer’s joint criterion as the anomaly score. We believe that this might put the Anomaly Transformer at some disadvantage in our experiments.. Following prior work (Goswami et al., 2023a), we downsample all time series longer than 2560 timesteps by a factor of 10 to speed up the training and evaluation process.
We report two anomaly detection metrics: adjusted best $F_{1}$ which is frequently used in practice (Goswami et al., 2023a; Challu et al., 2022), and the recently proposed volume under ROC surface (VUS-ROC) metric (Paparrizos et al., 2022a). For both metrics, higher scores are better.
Hyperparameters.
The hyperparameters used for training all models in our anomaly detection experiments are shown in Table 26.
Model Hyper-parameters
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial lr : $5e-5$
Anomaly Transformer sequence length : 512512512
number of channels : 111
k : 333
anomaly ratio : 4.004.004.00
model dimensions : 512512512
number of heads : 888
embedding layers : 333
dimension of feedforward layer : 512512512
DGHL sequence length : 512512512
number of channels : 111
hidden multiplier : 323232
max filters : 256256256
kernel multiplier : 111
sub-windows : 444
size of latent z vector : 505050
number of iteration in the Langevyn dynamics inference formula : 100100100
z step size : 0.10.10.1
noise std : 0.0010.0010.001
GPT4TS sequence length : 512512512
gpt layers : 333
patch length : 111
patch stride length : 111
transformer backbone : GPT-2
TimesNet sequence length : 512512512
dimension of model : 161616
dimension of feedforward layer : 161616
top k : 333
number of kernels : 666
$k$-NN k : 555
Table 26: Hyperparameter values for anomaly detection.
E.4 Imputation
Task Description.
Consider a time series $\mathcal{T}=[x_{1},...,x_{L}]$ and an observation mask $\mathcal{M}=[m_{1},...,m_{L}]$, where $m_{i}=$ 0 if $x_{i}$ is missing and $m_{i}=$ 1 if $x_{i}$ is observed. Then imputation is the task of estimating the missing values $\mathcal{T}$ by exploiting its observed values. We treat a patch as observed only if all its time steps are observed. For the remaining patches, we replace their patch embeddings with [MASK] and use MOMENT’s default reconstruction head to impute its values (Fig. 4 (iv)).
Datasets.
We evaluate imputation performance on 6 real-world datasets from domains where missing data is a common problem: 4 subsets of Electricity Transformer Temperature (ETT), Weather, and Electricity (Wu et al., 2023; Zhou et al., 2023).
Baselines.
We compare the two variants of MOMENT with 3 state-of-the-art deep learning methods, TimesNet, FPT, and DGHL; and 3 statistical interpolation methods, Cubic Spline, Linear, and 1-D Nearest Neighbor interpolation.
Experimental Setting.
To evaluate the models’ ability to interpolate missing values, we randomly mask contiguous sub-sequences of length 8. Instead of masking contiguous sub-sequences, previous studies (Wu et al., 2023; Zhou et al., 2023) mask individual time points, making the imputation task much easier. The results from prior studies are shown in Table 28. We observe that the statistical methods perform similarly to transformer methods, owing to the ease of the task. For our experiments involving randomly masking patches of length 8, our results are shown in Table 29. We measure the imputation performance of models using mean squared error, over 4 different masking rates: 12.5%, 25%, 37.5%, and 50%. f
Hyperparameters.
The hyperparameters used for training all models in our imputation experiments are shown in Table 27.
Model Hyper-parameters
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial lr : 0.00010.00010.0001
GPT4TS sequence length : 512512512
gpt layers : 333
patch length : 111
patch stride length : 111
transformer backbone : GPT-2
dimension of feedforward layer : 161616
TimesNet sequence length : 512512512
dimension of model : 646464
dimension of feedforward layer : 646464
top k : 333
number of kernels : 666
Table 27: Hyperparameter values for imputation.
Methods GPT4TS TimesNet PatchTST ETSformer LightTS DLinear FEDformer Stationary Autoformer Informer Reformer Naive Linear Nearest Cubic
Dataset Mask Ratio MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTm1$ 12.5% 0.017 0.085 0.023 0.101 0.041 0.130 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146 0.059 0.145 0.034 0.109 0.055 0.140 0.052 0.135
25% 0.022 0.096 0.023 0.101 0.044 0.135 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146 0.066 0.153 0.036 0.112 0.056 0.142 0.060 0.142
37.5% 0.029 0.111 0.029 0.111 0.049 0.143 0.133 0.271 0.113 0.231 0.103 0.219 0.069 0.191 0.039 0.131 0.057 0.161 0.079 0.200 0.063 0.182 0.077 0.164 0.038 0.117 0.060 0.146 0.071 0.151
50% 0.040 0.128 0.036 0.124 0.055 0.151 0.186 0.323 0.134 0.255 0.132 0.248 0.089 0.218 0.047 0.145 0.067 0.174 0.093 0.218 0.082 0.208 0.094 0.178 0.042 0.123 0.066 0.153 0.100 0.164
Avg 0.028 0.105 0.027 0.107 0.047 0.140 0.120 0.253 0.104 0.218 0.093 0.206 0.062 0.177 0.036 0.126 0.051 0.150 0.071 0.188 0.055 0.166 0.074 0.160 0.038 0.115 0.059 0.145 0.071 0.148
$ETTm2$ 12.5% 0.017 0.076 0.018 0.080 0.026 0.094 0.108 0.239 0.034 0.127 0.062 0.166 0.056 0.159 0.021 0.088 0.023 0.092 0.133 0.270 0.108 0.228 0.038 0.095 0.023 0.077 0.035 0.091 0.033 0.097
25% 0.020 0.080 0.020 0.085 0.028 0.099 0.164 0.294 0.042 0.143 0.085 0.196 0.080 0.195 0.024 0.096 0.026 0.101 0.135 0.272 0.136 0.262 0.041 0.100 0.025 0.081 0.036 0.093 0.039 0.103
37.5% 0.022 0.087 0.023 0.091 0.030 0.104 0.237 0.356 0.051 0.159 0.106 0.222 0.110 0.231 0.027 0.103 0.030 0.108 0.155 0.293 0.175 0.300 0.046 0.106 0.027 0.085 0.038 0.095 0.047 0.111
50% 0.025 0.095 0.026 0.098 0.034 0.110 0.323 0.421 0.059 0.174 0.131 0.247 0.156 0.276 0.030 0.108 0.035 0.119 0.200 0.333 0.211 0.329 0.051 0.115 0.030 0.090 0.041 0.100 0.062 0.122
Avg 0.021 0.084 0.022 0.088 0.029 0.102 0.208 0.327 0.046 0.151 0.096 0.208 0.101 0.215 0.026 0.099 0.029 0.105 0.156 0.292 0.157 0.280 0.044 0.104 0.026 0.083 0.038 0.095 0.045 0.109
$ETTh1$ 12.5% 0.043 0.140 0.057 0.159 0.093 0.201 0.126 0.263 0.240 0.345 0.151 0.267 0.070 0.190 0.060 0.165 0.074 0.182 0.114 0.234 0.074 0.194 0.211 0.275 0.083 0.183 0.181 0.260 0.107 0.207
25% 0.054 0.156 0.069 0.178 0.107 0.217 0.169 0.304 0.265 0.364 0.180 0.292 0.106 0.236 0.080 0.189 0.090 0.203 0.140 0.262 0.102 0.227 0.259 0.298 0.098 0.197 0.192 0.266 0.127 0.224
37.5% 0.072 0.180 0.084 0.196 0.120 0.230 0.220 0.347 0.296 0.382 0.215 0.318 0.124 0.258 0.102 0.212 0.109 0.222 0.174 0.293 0.135 0.261 0.323 0.326 0.119 0.215 0.215 0.277 0.160 0.245
50% 0.107 0.216 0.102 0.215 0.141 0.248 0.293 0.402 0.334 0.404 0.257 0.347 0.165 0.299 0.133 0.240 0.137 0.248 0.215 0.325 0.179 0.298 0.423 0.366 0.158 0.242 0.257 0.297 0.235 0.279
Avg 0.069 0.173 0.078 0.187 0.115 0.224 0.202 0.329 0.284 0.373 0.201 0.306 0.117 0.246 0.094 0.201 0.103 0.214 0.161 0.279 0.122 0.245 0.304 0.317 0.114 0.209 0.211 0.275 0.157 0.239
$ETTh2$ 12.5% 0.039 0.125 0.040 0.130 0.057 0.152 0.187 0.319 0.101 0.231 0.100 0.216 0.095 0.212 0.042 0.133 0.044 0.138 0.305 0.431 0.163 0.289 0.090 0.167 0.058 0.134 0.085 0.162 0.091 0.172
25% 0.044 0.135 0.046 0.141 0.061 0.158 0.279 0.390 0.115 0.246 0.127 0.247 0.137 0.258 0.049 0.147 0.050 0.149 0.322 0.444 0.206 0.331 0.097 0.176 0.060 0.138 0.088 0.165 0.101 0.179
37.5% 0.051 0.147 0.052 0.151 0.067 0.166 0.400 0.465 0.126 0.257 0.158 0.276 0.187 0.304 0.056 0.158 0.060 0.163 0.353 0.462 0.252 0.370 0.105 0.185 0.064 0.144 0.091 0.169 0.118 0.190
50% 0.059 0.158 0.060 0.162 0.073 0.174 0.602 0.572 0.136 0.268 0.183 0.299 0.232 0.341 0.065 0.170 0.068 0.173 0.369 0.472 0.316 0.419 0.118 0.199 0.071 0.153 0.097 0.176 0.150 0.205
Avg 0.048 0.141 0.049 0.146 0.065 0.163 0.367 0.436 0.119 0.250 0.142 0.259 0.163 0.279 0.053 0.152 0.055 0.156 0.337 0.452 0.234 0.352 0.102 0.182 0.063 0.142 0.090 0.168 0.115 0.186
$ECL$ 12.5% 0.080 0.194 0.085 0.202 0.055 0.160 0.196 0.321 0.102 0.229 0.092 0.214 0.107 0.237 0.093 0.210 0.089 0.210 0.218 0.326 0.190 0.308 0.214 0.293 0.079 0.182 0.181 0.271 0.091 0.196
25% 0.087 0.203 0.089 0.206 0.065 0.175 0.207 0.332 0.121 0.252 0.118 0.247 0.120 0.251 0.097 0.214 0.096 0.220 0.219 0.326 0.197 0.312 0.266 0.324 0.094 0.199 0.194 0.280 0.115 0.217
37.5% 0.094 0.211 0.094 0.213 0.076 0.189 0.219 0.344 0.141 0.273 0.144 0.276 0.136 0.266 0.102 0.220 0.104 0.229 0.222 0.328 0.203 0.315 0.339 0.366 0.117 0.223 0.220 0.296 0.152 0.245
50% 0.101 0.220 0.100 0.221 0.091 0.208 0.235 0.357 0.160 0.293 0.175 0.305 0.158 0.284 0.108 0.228 0.113 0.239 0.228 0.331 0.210 0.319 0.447 0.424 0.156 0.259 0.264 0.324 0.222 0.286
Avg 0.090 0.207 0.092 0.210 0.072 0.183 0.214 0.339 0.131 0.262 0.132 0.260 0.130 0.259 0.100 0.218 0.101 0.225 0.222 0.328 0.200 0.313 0.316 0.352 0.112 0.216 0.215 0.293 0.145 0.236
$Weather$ 12.5% 0.026 0.049 0.025 0.045 0.029 0.049 0.057 0.141 0.047 0.101 0.039 0.084 0.041 0.107 0.027 0.051 0.026 0.047 0.037 0.093 0.031 0.076 0.041 0.042 0.026 0.031 0.038 0.041 0.038 0.036
25% 0.028 0.052 0.029 0.052 0.031 0.053 0.065 0.155 0.052 0.111 0.048 0.103 0.064 0.163 0.029 0.056 0.030 0.054 0.042 0.100 0.035 0.082 0.045 0.045 0.027 0.032 0.040 0.041 0.043 0.039
37.5% 0.033 0.060 0.031 0.057 0.035 0.058 0.081 0.180 0.058 0.121 0.057 0.117 0.107 0.229 0.033 0.062 0.032 0.060 0.049 0.111 0.040 0.091 0.048 0.049 0.030 0.034 0.042 0.043 0.055 0.043
50% 0.037 0.065 0.034 0.062 0.038 0.063 0.102 0.207 0.065 0.133 0.066 0.134 0.183 0.312 0.037 0.068 0.037 0.067 0.053 0.114 0.046 0.099 0.054 0.054 0.033 0.037 0.044 0.045 0.069 0.048
Avg 0.031 0.056 0.030 0.054 0.060 0.144 0.076 0.171 0.055 0.117 0.052 0.110 0.099 0.203 0.032 0.059 0.031 0.057 0.045 0.104 0.038 0.087 0.047 0.048 0.029 0.034 0.041 0.043 0.051 0.042
Table 28: Results for the imputation task in the time-steps missing at random setting. Results averaged across 4 different masking rates: {12.5%, 25%, 37.5%, 50%}. Statistical interpolation methods such as forward and backward fill (naive), linear, nearest, and cubic interpolation perform better than many transformer-based baselines. Therefore, we consider the much harder, patches missing at random setting in our experiments.
Dataset Mask Ratio $\mathtt{MOMENT_{0}}$ $\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet Naive Linear Nearest Cubic
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 0.125 0.085 0.131 0.033 0.073 0.036 0.076 0.035 0.096 0.105 0.089 0.050 0.055 0.069 0.067 0.373 0.115
0.250 0.079 0.130 0.036 0.078 0.030 0.071 0.037 0.100 0.127 0.104 0.075 0.065 0.094 0.076 0.297 0.125
0.375 0.081 0.128 0.034 0.075 0.030 0.070 0.035 0.099 0.120 0.111 0.066 0.069 0.082 0.080 0.904 0.176
0.500 0.081 0.129 0.035 0.075 0.026 0.069 0.035 0.098 0.124 0.127 0.066 0.078 0.086 0.090 0.831 0.197
Mean 0.082 0.130 0.035 0.075 0.031 0.071 0.036 0.098 0.119 0.108 0.065 0.067 0.083 0.078 0.601 0.153
ETTh1 0.125 0.430 0.417 0.160 0.239 0.183 0.242 0.158 0.254 1.008 0.602 0.583 0.466 0.712 0.523 0.985 0.661
0.250 0.373 0.392 0.142 0.238 0.278 0.267 0.154 0.261 1.311 0.686 0.833 0.540 0.954 0.581 1.433 0.772
0.375 0.398 0.398 0.121 0.228 0.232 0.263 0.195 0.274 1.317 0.703 0.843 0.572 0.973 0.613 2.615 1.028
0.500 0.408 0.403 0.132 0.231 0.213 0.243 0.192 0.267 1.103 0.643 0.840 0.559 0.963 0.601 3.681 1.204
Mean 0.402 0.403 0.139 0.234 0.227 0.254 0.175 0.264 1.185 0.658 0.775 0.534 0.900 0.579 2.178 0.916
ETTh2 0.125 0.122 0.235 0.051 0.150 0.115 0.215 0.163 0.277 0.196 0.285 0.105 0.208 0.134 0.225 0.452 0.404
0.250 0.127 0.237 0.079 0.177 0.114 0.216 0.161 0.280 0.210 0.291 0.120 0.220 0.154 0.240 0.831 0.524
0.375 0.124 0.237 0.056 0.155 0.110 0.216 0.170 0.291 0.229 0.310 0.142 0.243 0.175 0.262 1.571 0.672
0.500 0.127 0.242 0.056 0.154 0.098 0.207 0.186 0.295 0.265 0.329 0.171 0.264 0.199 0.279 4.823 0.966
Mean 0.125 0.238 0.061 0.159 0.109 0.213 0.170 0.286 0.225 0.304 0.135 0.234 0.166 0.252 1.920 0.641
ETTm1 0.125 0.179 0.278 0.069 0.170 0.078 0.147 0.089 0.200 0.273 0.293 0.094 0.183 0.147 0.217 0.334 0.345
0.250 0.206 0.290 0.071 0.169 0.071 0.144 0.080 0.194 0.395 0.341 0.114 0.202 0.171 0.234 0.539 0.424
0.375 0.209 0.289 0.069 0.163 0.076 0.146 0.091 0.199 0.475 0.378 0.188 0.242 0.257 0.274 0.842 0.528
0.500 0.215 0.294 0.086 0.169 0.081 0.149 0.088 0.197 0.679 0.448 0.265 0.291 0.346 0.316 1.715 0.680
Mean 0.202 0.288 0.074 0.168 0.076 0.146 0.087 0.198 0.455 0.365 0.165 0.229 0.230 0.260 0.858 0.494
ETTm2 0.125 0.076 0.183 0.032 0.108 0.043 0.126 0.128 0.233 0.087 0.164 0.049 0.117 0.062 0.132 0.237 0.262
0.250 0.084 0.187 0.029 0.105 0.046 0.129 0.101 0.207 0.104 0.182 0.057 0.132 0.073 0.146 0.373 0.309
0.375 0.076 0.181 0.032 0.109 0.059 0.137 0.116 0.225 0.115 0.196 0.063 0.141 0.078 0.154 0.626 0.376
0.500 0.077 0.183 0.031 0.110 0.059 0.140 0.103 0.212 0.144 0.222 0.080 0.162 0.102 0.177 0.899 0.477
Mean 0.078 0.184 0.031 0.108 0.052 0.133 0.112 0.220 0.113 0.191 0.062 0.138 0.079 0.152 0.534 0.356
Electricity 0.125 0.251 0.370 0.095 0.211 0.069 0.180 0.126 0.248 1.350 0.818 0.458 0.466 0.608 0.492 0.924 0.610
0.250 0.249 0.372 0.093 0.211 0.073 0.184 0.121 0.246 1.447 0.857 0.654 0.554 0.815 0.582 1.619 0.769
0.375 0.250 0.371 0.094 0.211 0.071 0.181 0.125 0.248 1.518 0.888 0.836 0.637 1.031 0.675 2.507 0.959
0.500 0.250 0.371 0.092 0.210 0.075 0.185 0.126 0.249 1.581 0.915 1.002 0.712 1.239 0.766 3.978 1.213
Mean 0.250 0.371 0.094 0.211 0.072 0.183 0.124 0.248 1.474 0.869 0.737 0.592 0.923 0.629 2.257 0.888
Table 29: Imputation Results. MOMENT achieves state-of-the-art imputation results in both zero-shot and linear probe fine-tuning settings.
E.5 What is MOMENT Learning?
To investigate what MOMENT is learning, we conducted a series of experiments using synthetically generated sine waves to evaluate MOMENT’s ability to capture changes in trend, amplitude, frequencies, baselines, and phase of time series. In each experiment, $c$ controls the factor of interest, for example the power of the trend polynomial $c\in(\frac{1}{8},8)$ (Oreshkin et al., 2020) (Fig. 9), and frequency $c\in(1,32)$ of the generated sine waves (Fig. 9). We generate multiple sine waves by varying $c$, derive their sequence-level representations using MOMENT (Sec. 3.4), and visualize them in a 2- dimensional space using PCA and t-SNE (van der Maaten, 2014) in Fig. 4 and Fig. 7.
We also study the composition of the learnable mask embedding and the relationship between frequency and reconstruction error in a zero-shot setting. We find that the learned mask embedding is approximately composed of numbers drawn from the standard normal and that MOMENT can reconstruct lower frequency signals better. We observed a curious spike in reconstruction error around time series of frequency $c=64$. (Fig. 11)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(i) Trend (ii) Amplitude (iii) Frequency (iv) Basline Shift (v) Auto-correlation
Figure 7: What is MOMENT learning? Structure in the PCA (top) and t-SNE (bottom) visualizations of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and auto-correlation information. $\epsilon$ denotes gaussian noise with 0 mean and 0.10.10.1 standard deviation. $c$ controls the factor of interest, i.e. the power of the trend polynomial, amplitude, and frequency of the sine waves in experiments (i), (ii) & (iii), respectively.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: PCA (top) and t-SNE (bottom) visualizations of representations learned by MOMENT on the 5 largest UCR datasets. Different colors represent different classes. Even without dataset-specific fine-tuning, MOMENT learns distinct representations for different classes
(i) Trend
(ii) Amplitude
(iii) Frequency
(iv) Baseline Shift
(v) Phase
Figure 9: Examples of sinusoids used in the interpretability experiments.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Masking using a [MASK] tokens allows MOMENT to reconstruct time series in a zero-shot setting. Since zeros contain information, they bias model predictions. For two datasets ETTh1 and Weather, we mask the time series with zeros on the left and special mask tokens on the right.
Refer to caption Refer to caption
Figure 11: (Left) MOMENT can reconstruct lower frequency time series better in a zero-shot setting. (Right) The learned mask token is approximately composed of numbers drawn from a standard normal.
E.6 Impact of Model Size
We studied the impact of scaling the size of the model and training data on zero-shot forecasting, imputation, and anomaly detection performance. As shown in Fig. x, we found that increasing the size of the model generally improved zero-shot performance (lower MSE and sMAPE, higher VUS-ROC). Since varying the size of the pre-training dataset is expensive, we instead look at the zero-shot performance of model checkpoints before completing the first epoch. Our findings suggest that increasing the diversity in training data may also improve zero-shot performance.
Dataset Pred_horizon $\mathtt{MOMENT_{small}}$ $\mathtt{MOMENT_{base}}$ $\mathtt{MOMENT_{large}}$
MSE MAE MSE MAE MSE MAE
Weather 96 0.167 0.224 0.156 0.211 0.151 0.207
192 0.210 0.259 0.198 0.248 0.195 0.246
336 0.256 0.292 0.247 0.284 0.245 0.285
720 0.315 0.334 0.315 0.334 0.316 0.333
ETTh1 96 0.388 0.411 0.391 0.414 0.381 0.406
192 0.420 0.432 0.420 0.433 0.412 0.425
336 0.443 0.451 0.424 0.438 0.429 0.442
720 0.457 0.473 0.426 0.451 0.453 0.468
Table 30: Long-horizon forecasting scaling experiments for $\mathtt{MOMENT_{small}}$, $\mathtt{MOMENT_{base}}$, and $\mathtt{MOMENT_{large}}$.
Metric $\mathtt{MOMENT_{small}}$ $\mathtt{MOMENT_{base}}$ $\mathtt{MOMENT_{large}}$
Adj. F1 Mean 0.480 0.572 0.569
Median 0.450 0.641 0.607
Std. 0.378 0.383 0.372
Vus ROC Mean 0.643 0.677 0.660
Median 0.644 0.669 0.657
Std. 0.137 0.121 0.130
Table 31: Zero-shot anomaly detection scaling experiments for $\mathtt{MOMENT_{small}}$, $\mathtt{MOMENT_{base}}$, and $\mathtt{MOMENT_{large}}$.
$\mathtt{MOMENT_{small}}$ $\mathtt{MOMENT_{base}}$ $\mathtt{MOMENT_{large}}$
Mean 0.716705 0.766437 0.763933
Median 0.720930 0.771078 0.766667
Std. 0.155945 0.156763 0.160345
Avg. rank 2.686813 1.659341 1.653846
Median rank 3.000000 2.000000 1.50
Wins/Losses 28.5/153.5 122.0/60.0 122.5/59.5
Table 32: Zero-shot classification scaling experiments for $\mathtt{MOMENT_{small}}$, $\mathtt{MOMENT_{base}}$, and $\mathtt{MOMENT_{large}}$.
E.7 Training losses
Refer to caption Refer to caption Refer to caption
Figure 12: Training losses (MSE). A dashed vertical line denotes the first epoch. All models were trained with a batch size of 131072 patches. (left) Larger models obtain lower training loss. right Eventually, randomly initialized MOMENT-small outperform the same model initialized with Flan-T5 weights. The figure on the right is in log scale.
E.8 Efficiency Analysis
Model ETTh1-96
Total Param. (M) Trainable Param. (M) Mem. (MiB)
MOMENT 347.53 6.29 2079
GPT4TS 82.28 1.12 1031
TimesNet 0.89 0.89 683
Time-LLM 3623.71 254.37 4537
Table 33: Efficiency analysis of MOMENT against other forecasting models on the ETTh1 with prediction horizon set to 96. MOMENT outperforms all the listed models and has a fraction of parameters as the most recent LLM-based forecasting method.
Appendix F Transparency Index
Sub-domain Indicator MOMENT
Data Data size 1
Data sources 1
Data creators 0
Data source selection 1
Data curation 1
Data augmentation 1
Harmful data filtration 0
Copyrighted data 1
Data license 1
Personal information in data 1
Data Labor Use of human labor 1
Employment of data laborers 1
Geographic distribution of data laborers 1
Wages 1
Instructions for creating data 1
Labor protections 1
Third party partners 1
Data Access Queryable external data access 1
Direct external data access 1
Compute Compute usage 1
Development duration 1
Compute hardware 1
Hardware owner 1
Energy usage 1
Carbon emissions 1
Broader environmental impact 0
Methods Model stages 1
Model objectives 1
Core frameworks 1
Additional dependencies 1
Data Mitigations Mitigations for privacy 0
Mitigations for copyright 0
Upstream Subtotal 75%
Sub-domain Indicator MOMENT
Model Basics Input modality 1
Output modality 1
Model components 1
Model size 1
Model architecture 1
Centralized model documentation 1
Model Access External model access protocol 1
Blackbox external model access 1
Full external model access 1
Capabilities Capabilities description 1
Capabilities demonstration 1
Evaluation of capabilities 1
External reproducibility of capabilities evaluation 0
Third party capabilities evaluation 0
Limitations Limitations description 1
Limitations demonstration 1
Third party evaluation of limitations 0
Risks Risks description 1
Risks demonstration 0
Unintentional harm evaluation 0
External reproducibility of unintentional harm evaluation 0
Intentional harm evaluation 0
External reproducibility of intentional harm evaluation 0
Third party risks evaluation 0
Model Mitigations Mitigations description 0
Mitigations demonstration 0
Mitigations Mitigations evaluation 0
External reproducibility of mitigations evaluation 0
Third party mitigations evaluation 0
Trustworthiness Trustworthiness evaluation 0
External reproducibility of trustworthiness evaluation 0
Inference Inference duration evaluation 1
Inference compute evaluation 1
Model Subtotal 51.5%
Table 34: Expected (left) upstream and (right) model transparency scores. MOMENT has one of the highest upstream transparency. Our model transparency scores are lower due to (third-party) harm, mitigations, trustworthiness evaluation, which are not well understood for time series modeling.
Appendix G Results Sources
Task Method Type Reimplementation/ Rerun Source
Long-horizon Forecasting Time-LLM LLM-based $\checkmark$ Time-LLM
GPT4TS $\times$ One Fits All
PatchTST, Fedformer, Autoformer,
Stationary, ETSformer, LightTS,
Informer, Reformer
Transformer-based $\times$ One Fits All
Pyraformer, LogTrans $\times$ TimesNet
TimesNet, DLinear Deep learning $\times$ One Fits All
N-BEATS $\checkmark$ N-BEATS
Short-horizon Forecasting GPT4TS LLM-based $\checkmark$ One Fits All
TimesNet Deep learning $\checkmark$ TimesNet
N-BEATS $\checkmark$ N-BEATS
AutoARIMA, AutoTheta, AutoETS,
Seasonal Naive, Naive, Random Walk
Statistical learning $\checkmark$ Nixtla Statsforecast Repository
Classification GPT4TS LLM-based $\checkmark$ One Fits All
TimesNet Deep learning $\checkmark$ TimesNet
TS2Vec, T-Loss, TNC, TS-TCC, TST Unsupervised Representation learning $\times$ TS2Vec
CNN, Encoder, FCN, MCNN, MLP,
ResNet, t-LeNet, TWIESN
Deep learning $\times$ DL4TSC Repository
DTW Statistical learning $\times$ TS2Vec
Anomaly Detection GPT4TS LLM-based $\checkmark$ One Fits All
TimesNet Deep learning $\checkmark$ TimesNet
Anomaly Transformer Transformer-based $\checkmark$ Anomaly Transformer
DGHL Deep learning $\checkmark$ Time Series Model Selection
$k$-NN Statistical learning $\checkmark$ Time Series Model Selection
Imputation GPT4TS LLM-based $\checkmark$ One Fits All
TimesNet Deep learning $\checkmark$ TimesNet
PatchTST, ETSformer, LightTS,
Fedformer,Stationary, Autoformer,
Informer, Reformer
Transformer-based $\times$ One Fits All
DLinear Deep learning $\times$ One Fits All
Naive Statistical learning $\checkmark$ Pandas FFill, Pandas BFill
Linear, Nearest, Cubic $\checkmark$ Scipy Interp1D
Table 35: Source for the results for each baseline for all downstream task.
Appendix H Radar Plot
We generate a radar plot (Fig. 1) to visually compare MOMENT with GPT4TS and TimesNet. The values obtained by each method for a given task are min-max normalized with respect to the other methods for each of the 5 downstream tasks. For imputation, long- and short-horizon forecasting, we report 11-1 - the normalized MSE or sMAPE for the methods on the weather and (subset of) M4 datasets, respectively. For classification and anomaly detection, we report the average accuracy and VUS-ROC of the methods across all the datasets.

### v3

@@
 1 Introduction
-Figure 1: MOMENT can solve multiple time series analysis tasks well (App. E).
+Figure 1: MOMENT can solve multiple time series analysis tasks well (App. F).
 Time series analysis is an important field encompassing a wide range of applications ranging from forecasting weather patterns (Schneider & Dickinson, 1974) or detecting irregular heartbeats using Electrocardiograms (Goswami et al., 2021), to identifying anomalous software deployments (Xu et al., 2018).
@@
 (2020) to the original Transformer (Vaswani et al., 2017).
-Specifically, we remove the additive bias from the Layer Norm (Ba et al., 2016), and place it before the residual skip connections (He et al., 2016), and use the relation positional embedding scheme (Shaw et al., 2018).
+Specifically, we remove the additive bias from the Layer Norm (Ba et al., 2016), and place it before the residual skip connections (He et al., 2016), and use the relative positional embedding scheme (Shaw et al., 2018).
 Below we summarize the intuition behind some of our key design decisions.
@@
 We use a lightweight prediction head instead of a decoder of the same size as the encoder, to enable the necessary architectural modifications for task-specific fine-tuning of a limited number of trainable parameters while keeping the bulk of parameters and the high-level features learned by the encoder intact.
+Additional absolute positional embeddings.
+In addition to relative positional embeddings, we add absolute sinusoidal positional embeddings (Vaswani et al., 2017) to each patch [^3] We initially included absolute positional embeddings by accident, but subsequent experiments showed that combining absolute and relative positional embeddings improve predictions..
 Tasks Supervision Datasets Metrics Baselines Experimental Setting
@@
 Look-back window $L=512$,
-[…]
+Forecast horizon $H=\{24,60\}$ (ILI), $\{96,720\}$ (rest)
 Short-horizon
@@
 UCR Classification
-Archive (subset)
+Archive
 Accuracy
@@
 UCR Anomaly
-Archive (subset)
+Archive
 Adjusted Best F1
@@
 Anomaly Nearest Neighbor
-Reconstruction-based anomaly detection with window size =512=512= 512
+Reconstruction-based anomaly detection with window size $=512$
 MSE between observed and predicted time series
@@
 length 8
-[…]
+Masking ratios: $\{12.5\%,25\%,37.5\%,50\%\}$
 Table 1: Experimental benchmark. We evaluate MOMENT on 5 time series analysis tasks with an emphasis on limited memory, compute, and supervision settings.
@@
 We use gradient checkpointing (Radford et al., 2021) to improve training throughput and save memory, and train all models in a mixed precision setting, using float-32 for numerically unstable operations, e.g.
-layer normalization, and bfloat-16444https://cloud.google.com/tpu/docs/bfloat16, otherwise.
+layer normalization, and bfloat-16555https://cloud.google.com/tpu/docs/bfloat16, otherwise.
 We train all models for 2 epochs.
@@
 We provide detailed descriptions of each task and MOMENT’s configuration in App.
-E.
+F.
 Fine-tuning settings.
@@
 18.
-$\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet N-BEATS ARIMA Theta ETS
-Seasonal
-Naive
-Naive
-Random
-Walk
+$\mathtt{MOMENT_{LP}}$ GPT4TS TimesNet N-BEATS ARIMA Theta ETS Seasonal Naive Naive Random Walk
 Datasets M4 FR M4 FR M4 FR M4 FR
@@
 As is common in prior work (Yue et al., 2022; Franceschi et al., 2019), the quality of representations is measured using the accuracy of a Support Vector Machine trained on them (App.
-E.2).
+F.2).
 For short-horizon forecasting, we consider the zero-shot setting introduced by Oreshkin et al.
 (2021).
-In particular, we fine-tune MOMENT on a source dataset using a forecasting head, and evaluate its performance on a target dataset without any fine-tuning (App E.1.2, Tab.
+In particular, we fine-tune MOMENT on a source dataset using a forecasting head, and evaluate its performance on a target dataset without any fine-tuning (App F.1.2, Tab.
 21).
@@
 1 and detailed in App.
-E.
+F.
 Metrics.
@@
 Hyper-parameter tuning.
-[…]
+We do not perform hyper-parameter tuning.
+In all experiments that follow, unless mentioned otherwise, we fine-tune MOMENT-Large with a batch size of 64, and one cycle learning rate schedule with a peak learning rate between $55e-$ 5 and $31e-$ 3 (Smith & Topin, 2019).
+For baseline methods, we capture recommended settings from their papers and public repositories.
+We report all hyper-parameters settings for MOMENT and baselines in App.
+F.
 Research questions. Through the following experiments we aim to answer 3 broad research questions.
@@
 Std. 0.147 0.238 0.234 0.134 0.136 0.168 0.176 0.220 0.180 0.159 0.188 0.194 0.169 0.177 0.221 0.164 0.152
-Mean rank 7.225 13.324 13.318 3.494 5.261 6.937 6.500 11.846 9.384 8.906 5.565 11.043 9.247 4.406 16.115 10.384 9.071
+Mean rank 7.2 13.3 13.3 3.5 5.3 6.9 6.5 11.8 9.4 8.9 5.6 11.0 9.2 4.4 16.1 10.4 9.1
 Median rank 7.0 14.0 14.0 3.0 5.0 6.5 6.0 13.0 9.0 9.0 4.0 12.0 10.0 3.0 17.0 11.0 9.0
-Wins/Losses 880.5/566.5 325.5/1121.5 326.0/1121.0 1220.0/227.0 1033.0/375.0 885.5/522.5 924.0/484.0 460.0/987.0 684.0/763.0 727.5/719.5 1031.5/415.5 533.0/914.0 696.5/750.5 1137.0/310.0 71.5/1375.5 593.0/854.0 712.5/734.5
+Wins 880 325 326 1220 1033 885 924 460 684 727 1031 533 696 1137 71 593 712
+Losses 566 1121 1121 227 375 522 484 987 763 719 415 914 750 310 1375 854 734
 Table 4: Classification accuracy of methods across 91 UCR datasets.
@@
 23.
+Results on multi-variate UEA datasets are in Tab.
+24.
 4.1 MOMENT can solve multiple time series modeling tasks in limited supervision settings
@@
 Median rank 3.00 3.00 5.00 3.50 4.00 5.25 3.75
-Wins/Losses 703.5/472.5 783.0/393.0 419.0/757.0 658.0/518.0 524.0/652.0 378.0/798.0 650.5/525.5
+Wins/Losses 703/472 783/393 419/757 658/518 524/652 378/798 650/525
 VUS ROC Mean 0.670 0.684 0.611 0.679 0.661 0.646 0.706
@@
 Median rank 4.00 3.00 6.00 4.00 4.00 4.50 3.00
-Wins/Losses 577.0/599.0 709.0/467.0 354.0/822.0 608.0/568.0 605.0/571.0 509.0/667.0 754.0/422.0
+Wins/Losses 577/599 709/467 354/822 608/568 605/571 509/667 754/422
 Table 7: Anomaly detection performance averaged over 248 time series from the UCR Anomaly Archive.
@@
 Lastly, we would like to thank Cecilia Morales for her insightful comments, especially on the broader impacts of this work, and for helping us proofread this manuscript.
+Finally, we would like to thank the anonymous ICML reviewers for their comments which helped improve our work.
 Data.
@@
 We are working on open-sourcing MOMENT-base and MOMENT-small, and our research code public.
-The latter is currently available anonymously at https://anonymous.4open.science/r/BETT-773F/README.md.
+The latter is currently available at https://github.com/moment-timeseries-foundation-model/moment-research.
 We enlist an exhaustive list of hyper-parameters in App.
-E to aid reproducibility.
+F to aid reproducibility.
 We would like to emphasize that all datasets used in this study are publicly available.
@@
 Ethical considerations and potential misuse.
-Despite MOMENT’s promising performance in limited-data settings, it is important to use its predictions with care, especially in high-stakes settings such as healthcare.
-Before MOMENT is used for high-stakes decision-making, we recommend fine-tuning and evaluating the model with task-specific in-domain data.
-Appendix A Related Work
+Despite MOMENT’s promising performance in limited-data settings, it’s important to use its predictions with care, especially in high-stakes settings like healthcare.
+The reliability of MOMENT’s predictions hinges on the quality and diversity of the data it was trained on.
+Any dependence on unreliable or biased data could potentially result in skewed outputs, unfairly portraying or discriminating against certain groups or individuals.
+While MOMENT demonstrates the ability to capture changes time-series characteristics, e.g.
+trend, amplitude, frequency, it struggles to differentiate between vertically shifted time-series as it normalizes signals prior to modeling.
+Greater interpretability and explainability of MOMENT’s predictions are necessary to enable more informed decision-making processes.
+Prior to utilizing MOMENT in high-stakes decision-making, we recommend fine-tuning and evaluating the model with task-specific, in-domain data to address these ethical concerns and ensure equitable outcomes.
+Additionally, efforts should be made to promote equitable access to MOMENT and other AI technologies, particularly within healthcare systems where disparities in access may exacerbate existing inequities.
+Appendix A Changelog
+  • (Oct’2024) Added links to both research and inference benchmarks, notes on contemporary work in Tab.
+9 (Timer (Liu et al., 2024) and TOTEM (Talukder et al., 2024)).
+We also note differences in positional embeddings (Sec.
+3.2).
+Appendix B Related Work
 Transformers and Patching for Time series Modeling.
@@
 In our work, we consider both these questions and show that pre-training a model of sufficient capacity on a large corpus of unlabeled time series data can in fact enable it to provide reasonably accurate predictions in limited-supervision settings.
-Appendix B Contemporary Work
-Feature MOMENT Moirai Lag-LLaMa Chronos TimesFM TimeGPT-1 TinyTimeMixers TOTEM
-Citation Ours (Woo et al., 2024) (Rasul et al., 2023) (Ansari et al., 2024) (Das et al., 2023) (Garza & Mergenthaler-Canseco, 2023) (Ekambaram et al., 2024) (Talukder et al., 2024)
-Base Architecture T5 encoder Encoder-only transformer LLaMa T5 (encoder-decoder) Decoder-only Transformer MLP-Mixer (Tolstikhin et al., 2021) VQ-VAE (Van Den Oord et al., 2017)
-Sizes 40M, 125M, 385M 14M, 91M, 311M 200M 20M, 46M, 200M, 710M 17M, 70M, 200M ? $<$1M 0.3M
-Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Appendix C Contemporary Work
+Feature MOMENT Moirai Lag-LLaMa Chronos Timer
+Citation Ours (Woo et al., 2024) (Rasul et al., 2023) (Ansari et al., 2024) (Liu et al., 2024)
+Base Architecture T5 encoder Encoder-only transformer LLaMa T5 (encoder-decoder) Decoder-only
+Sizes 40M, 125M, 385M 14M, 91M, 311M 200M 20M, 46M, 200M, 710M 29M, 50M, 67M
+Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Evaluation Tasks
@@
 Anomaly detection, Imputation
-Forecasting Forecasting Forecasting Forecasting Forecasting Forecasting
-Forecasting, Classification,
-Anomaly detection, Imputation
-Data Total Observations 1.13B++ 27.65B [^9] These values are taken from Tables 8–17 of (Woo et al., 2024) and denote the # of observations 0.36B [^10] These values are taken from Table 4 of (Rasul et al., 2023) and denote the # of observations 84B 100B 100B 244M N/A
-# Time Series 13M 4M 11K 11M ? ? N/A N/A
-Public $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Public Release Upon acceptance Feb’2024 Feb’2024 Mar’2024 - - Apr’2024 Feb’2024
-Tokenization Fixed-length patches Multi-scale patches Lag features Scaling & quantization Fixed-length patches ? Adaptive Patching VQVAE-based
+Forecasting Forecasting Forecasting
+Forecasting, Imputation,
+Anomaly detection
+Data Total Observations 1.13B++ 27.65B [^10] These values are taken from Tables 8–17 of (Woo et al., 2024) and denote the # of observations 0.36B [^11] These values are taken from Table 4 of (Rasul et al., 2023) and denote the # of observations 84B 28B
+# Time Series 13M 4M 11K 11M 7.4M
+Public $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Public Release May’24 Feb’2024 Feb’2024 Mar’2024 Jun’2024
+Tokenization Fixed-length patches Multi-scale patches Lag features Scaling & quantization Single-series Sequence
 Objective
@@
 of Student’s t distribution
-Forecast Cross-entropy loss Forecasting error (MSE) ? Forecasting error (MSE)
+Forecast Cross-entropy loss Next token prediction
+Distributional Prediction In progress (post hoc) $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+Interpretability $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+Any-variate Prediction $\checkmark$ (Channel independence) $\checkmark$ (Flattening) $\times$ $\times$ $\checkmark$ (Uniform sub-sampling)
+Context length 512 1000–5000 1024 512 1440
+Scaling Model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Data $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
+Impact of Initialization $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+Feature TimesFM TimeGPT-1 TinyTimeMixers TOTEM
+Citation (Das et al., 2023) (Garza & Mergenthaler-Canseco, 2023) (Ekambaram et al., 2024) (Talukder et al., 2024)
+Base Architecture Decoder-only Transformer MLP-Mixer (Tolstikhin et al., 2021) VQ-VAE (Van Den Oord et al., 2017)
+Sizes 17M, 70M, 200M ? $<$1M 0.3M
+Open Source $\times$ $\times$ $\checkmark$ $\checkmark$
+Evaluation Tasks Forecasting Forecasting Forecasting
+Forecasting, Classification,
+Anomaly detection, Imputation
+Data Total Observations 100B 100B 244M N/A
+# Time Series ? ? N/A N/A
+Public $\times$ $\times$ $\checkmark$ $\checkmark$
+Public Release - - Apr’2024 Feb’2024
+Tokenization Fixed-length patches ? Adaptive Patching VQVAE-based
+Objective Forecasting error (MSE) ? Forecasting error (MSE)
 Reconstruction +
 Commitment error
-Distributional Prediction In progress (post hoc) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ (post hoc) $\times$ $\checkmark$
-Interpretability $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$
-Any-variate Prediction $\checkmark$ (Channel independence) $\checkmark$ (Flattening) $\times$ $\times$ $\times$ $\checkmark$ (?) $\checkmark$ (Channel Independence + Mixing) $\checkmark$ (Flattening)
-Context length 512 1000–5000 1024 512 512 ? 512 512
-Scaling Model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
-Data $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
-Impact of Initialization $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+Distributional Prediction $\times$ $\checkmark$ (post hoc) $\times$ $\checkmark$
+Interpretability $\checkmark$ $\times$ $\times$ $\times$
+Any-variate Prediction $\times$ $\checkmark$ (?) $\checkmark$ (Channel Independence + Mixing) $\checkmark$ (Flattening)
+Context length 512 ? 512 512
+Scaling Model $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Data $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Impact of Initialization $\times$ $\times$ $\checkmark$ $\times$
 Table 9: Time series foundation modeling has a young but growing literature.
@@
 Moirai (Woo et al., 2024) uses a mixture of Student’s t, log-normal, negative binomial, and low variance normal distribution.
-Appendix C Interesting directions for future work
+Appendix D Interesting directions for future work
 We note some interesting directions for future work:
-  • Study the impact of design choices such as the impact of the choice of the loss function (Huber, $L_{1}$, $L_{2}$), patch length (4,8)(4,8)( 4 , 8 ), and masking percentage (0.3,0.6)(0.3,0.6)( 0.3 , 0.6 ) on pre-training loss and time series modeling performance.
+  • Study the impact of design choices such as the impact of the choice of the loss function (Huber, $L_{1}$, $L_{2}$), patch length $(4,8)$, and masking percentage $(0.3,0.6)$ on pre-training loss and time series modeling performance.
   • Pre-training data.
 Two interesting directions include using augmentation and synthetic data (Ansari et al., 2024) to improve the quality of pre-training, and looking at tuning dataset mixtures in the Time Series Pile.
-Appendix D The Time Series Pile
+Appendix E The Time Series Pile
 We compiled a large collection of publicly available datasets from diverse domains into the Time Series Pile.
@@
 Task Dataset Channels Series Length Data Size (Train, Val, Test) Information (Frequency/Number of Classes)
-Long
-horizon
-forecasting
-ETTm1, ETTm2 7 {96, 720} (33953, 11425, 11425) Electricity (15 mins)
+Long horizon forecasting (Informer) ETTm1, ETTm2 7 {96, 720} (33953, 11425, 11425) Electricity (15 mins)
 ETTh1, ETTh2 7 (8033, 2785, 2785) Electricity (15 mins)
@@
 ILI 7 {24, 60} (69, 2, 98) Illness (Weekly)
-Short
-horizon
-forecasting
-M4-Yearly 1 6 (16099, 2301, 4600) -
+Short horizon forecasting (Monash) M4-Yearly 1 6 (16099, 2301, 4600) -
 M4-Quarterly 8 (16800, 2400, 4800) -
@@
 M3-Monthly 18 (999, 144, 285) -
-Imputation
-ETTm1, ETTm2 7 512 (33953, 11425, 11425) Electricity (15 mins)
+Imputation (Informer) ETTm1, ETTm2 7 512 (33953, 11425, 11425) Electricity (15 mins)
 ETTh1, ETTh2 7 (8033, 2785, 2785) Electricity (15 mins)
@@
 Weather 21 (36280, 5175, 10444) Weather (10 mins)
-Classification
-UWaveGestureLibraryX 1 315 (640, 256, 3582) Motion Gesture (8 classes)
+Classification (UCR) UWaveGestureLibraryX 1 315 (640, 256, 3582) Motion Gesture (8 classes)
 ECG5000 140 (357, 143, 4500) ECG Record (5 classes)
@@
 Ham 431 (77, 32, 105) Food spectrographs (2 classes)
-Anomaly
-detection
-(TSB-UAD)
-1sddb40 1 - (24489, 9489, 3969) Beats
+Anomaly detection (TSB-UAD) 1sddb40 1 - (24489, 9489, 3969) Beats
 BIDMC1 - (1274, 204, 7988) PVC
@@
 Healthcare200 Healthcare 200 2 1 96 200 19200
-Appendix E Experimental Setup and Results
+Appendix F Experimental Setup and Results
 Through our experiments, our goal is to answer the following research questions.
@@
 We report all hyper-parameters settings for MOMENT and baselines in Appendix D.
-E.1 Forecasting
+F.1 Forecasting
 Task description.
@@
 4 (i)).
-E.1.1 Long-Horizon Forecasting
+F.1.1 Long-Horizon Forecasting
 Datasets.
@@
 Model Hyper-parameters
-MOMENT sequence length : 512512512
-patch length : 888
-patch stride length : 888
-initial learning rate : 0.00010.00010.0001
-[…]
-Time-LLM sequence length : 512512512
-patch length : 161616
-patch stride length : 888
-initial learning rate : 0.0010.0010.001
-dimension of feedforward layer : 204820482048
-llm layers : 323232
-number of heads : 888
-N-BEATS sequence length : 512512512
+MOMENT sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+initial learning rate : $0.0001$
+forecast horizon : $\{96,720\}$
+Time-LLM sequence length : $512$
+patch length : $16$
+patch stride length : $8$
+initial learning rate : $0.001$
+dimension of feedforward layer : $2048$
+llm layers : $32$
+number of heads : $8$
+N-BEATS sequence length : $512$
 stack types : {trend, seasonality}
-number of blocks per stack : 333
-[…]
-hidden layer units : 256256256
+number of blocks per stack : $3$
+thetas dimensions : $\{4,8\}$
+hidden layer units : $256$
 Table 17: Hyper-parameter values for long-horizon forecasting models.
@@
 Table 19: Long-term forecasting performance with a look-back window of 96 time-steps. Results are taken from iTransformer (Liu et al., 2023).
-E.1.2 Zero-shot Short-Horizon Forecasting
+F.1.2 Zero-shot Short-Horizon Forecasting
 Datasets.
@@
 Model Hyper-parameters
-$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-initial learning rate : 0.0020.0020.002
-[…]
-$\mathtt{MOMENT_{0}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-initial learning rate : 0.0010.0010.001
-N-BEATS sequence length : 512512512
+$\mathtt{MOMENT_{LP}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+initial learning rate : $0.002$
+max epochs : $\{5,10\}$
+$\mathtt{MOMENT_{0}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+initial learning rate : $0.001$
+N-BEATS sequence length : $512$
 stack types : {’trend’, ’seasonality’}
-number of blocks per stack : 333
-[…]
-hidden layer units : 256256256
-GPT4TS forecast horizon : 0
-gpt layers : 333
-patch length : 111
-patch stride length : 111
-sequence length : 512512512
-TimesNet sequence length : 512512512
-model dimension : 323232
-dimension of feedforward layer : 323232
-[…]
+number of blocks per stack : $3$
+thetas dimensions : $\{4,8\}$
+hidden layer units : $256$
+GPT4TS forecast horizon : 00
+gpt layers : $3$
+patch length : $1$
+patch stride length : $1$
+sequence length : $512$
+TimesNet sequence length : $512$
+model dimension : $32$
+dimension of feedforward layer : $32$
+top-$k$ : $5$
 Table 20: Hyper-parameter values for short-horizon forecasting models.
@@
 Table 24: Classification accuracy of methods across 29 UEA datasets. MOMENT without fine-tuning on individual datasets demonstrates promising accuracy.
-E.2 Classification
+F.2 Classification
 Task Description.
@@
 Model Hyper-parameters
-$\mathtt{MOMENT_{0}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-[…]
+$\mathtt{MOMENT_{0}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+$\mathtt{SVM}$ C : $\{0.0001,0.001,0.01,0.1,1,10,100,1000,10000\}$
 kernel : RBF
-degree : 333
-cache size : 200200200
-max iterations : 100000001000000010000000
+degree : $3$
+cache size : $200$
+max iterations : $10000000$
 decision function shape : One versus rest
 Table 25: Hyper-parameter values for classification.
-E.3 Anomaly Detection
+F.3 Anomaly Detection
 Task Description.
@@
 Model Hyper-parameters
-$\mathtt{MOMENT_{0}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-initial lr : $5e-5$
-Anomaly Transformer sequence length : 512512512
-number of channels : 111
-k : 333
-anomaly ratio : 4.004.004.00
-model dimensions : 512512512
-number of heads : 888
-embedding layers : 333
-dimension of feedforward layer : 512512512
-DGHL sequence length : 512512512
-number of channels : 111
-hidden multiplier : 323232
-max filters : 256256256
-kernel multiplier : 111
-sub-windows : 444
-size of latent z vector : 505050
-number of iteration in the Langevyn dynamics inference formula : 100100100
-z step size : 0.10.10.1
-noise std : 0.0010.0010.001
-GPT4TS sequence length : 512512512
-gpt layers : 333
-patch length : 111
-patch stride length : 111
+$\mathtt{MOMENT_{0}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+$\mathtt{MOMENT_{LP}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+initial lr : $55e-$ 5
+Anomaly Transformer sequence length : $512$
+number of channels : $1$
+k : $3$
+anomaly ratio : $4.00$
+model dimensions : $512$
+number of heads : $8$
+embedding layers : $3$
+dimension of feedforward layer : $512$
+DGHL sequence length : $512$
+number of channels : $1$
+hidden multiplier : $32$
+max filters : $256$
+kernel multiplier : $1$
+sub-windows : $4$
+size of latent z vector : $50$
+number of iteration in the Langevyn dynamics inference formula : $100$
+z step size : $0.1$
+noise std : $0.001$
+GPT4TS sequence length : $512$
+gpt layers : $3$
+patch length : $1$
+patch stride length : $1$
 transformer backbone : GPT-2
-TimesNet sequence length : 512512512
-dimension of model : 161616
-dimension of feedforward layer : 161616
-top k : 333
-number of kernels : 666
-$k$-NN k : 555
+TimesNet sequence length : $512$
+dimension of model : $16$
+dimension of feedforward layer : $16$
+top k : $3$
+number of kernels : $6$
+$k$-NN k : $5$
 Table 26: Hyperparameter values for anomaly detection.
-E.4 Imputation
+F.4 Imputation
 Task Description.
@@
 Model Hyper-parameters
-$\mathtt{MOMENT_{0}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
-patch length : 888
-patch stride length : 888
-initial lr : 0.00010.00010.0001
-GPT4TS sequence length : 512512512
-gpt layers : 333
-patch length : 111
-patch stride length : 111
+$\mathtt{MOMENT_{0}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+$\mathtt{MOMENT_{LP}}$ sequence length : $512$
+patch length : $8$
+patch stride length : $8$
+initial lr : $0.0001$
+GPT4TS sequence length : $512$
+gpt layers : $3$
+patch length : $1$
+patch stride length : $1$
 transformer backbone : GPT-2
-dimension of feedforward layer : 161616
-TimesNet sequence length : 512512512
-dimension of model : 646464
-dimension of feedforward layer : 646464
-top k : 333
-number of kernels : 666
+dimension of feedforward layer : $16$
+TimesNet sequence length : $512$
+dimension of model : $64$
+dimension of feedforward layer : $64$
+top k : $3$
+number of kernels : $6$
 Table 27: Hyperparameter values for imputation.
@@
 Table 29: Imputation Results. MOMENT achieves state-of-the-art imputation results in both zero-shot and linear probe fine-tuning settings.
-E.5 What is MOMENT Learning?
+F.5 What is MOMENT Learning?
 To investigate what MOMENT is learning, we conducted a series of experiments using synthetically generated sine waves to evaluate MOMENT’s ability to capture changes in trend, amplitude, frequencies, baselines, and phase of time series.
@@
 Structure in the PCA (top) and t-SNE (bottom) visualizations of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and auto-correlation information.
-$\epsilon$ denotes gaussian noise with 0 mean and 0.10.10.1 standard deviation.
+$\epsilon$ denotes gaussian noise with 00 mean and $0.1$ standard deviation.
 $c$ controls the factor of interest, i.e.
@@
 Figure 11: (Left) MOMENT can reconstruct lower frequency time series better in a zero-shot setting. (Right) The learned mask token is approximately composed of numbers drawn from a standard normal.
-E.6 Impact of Model Size
+F.6 Impact of Model Size
 We studied the impact of scaling the size of the model and training data on zero-shot forecasting, imputation, and anomaly detection performance.
@@
 Table 32: Zero-shot classification scaling experiments for $\mathtt{MOMENT_{small}}$, $\mathtt{MOMENT_{base}}$, and $\mathtt{MOMENT_{large}}$.
-E.7 Training losses
+F.7 Training losses
 Refer to caption Refer to caption Refer to caption
@@
 The figure on the right is in log scale.
-E.8 Efficiency Analysis
+F.8 Efficiency Analysis
 Model ETTh1-96
@@
 MOMENT outperforms all the listed models and has a fraction of parameters as the most recent LLM-based forecasting method.
-Appendix F Transparency Index
+Appendix G Transparency Index
 Sub-domain Indicator MOMENT
@@
 Our model transparency scores are lower due to (third-party) harm, mitigations, trustworthiness evaluation, which are not well understood for time series modeling.
-Appendix G Results Sources
+Appendix H Results Sources
 Task Method Type Reimplementation/ Rerun Source
@@
 Table 35: Source for the results for each baseline for all downstream task.
-Appendix H Radar Plot
+Appendix I Radar Plot
 We generate a radar plot (Fig.
@@
 The values obtained by each method for a given task are min-max normalized with respect to the other methods for each of the 5 downstream tasks.
-For imputation, long- and short-horizon forecasting, we report 11-1 - the normalized MSE or sMAPE for the methods on the weather and (subset of) M4 datasets, respectively.
+For imputation, long- and short-horizon forecasting, we report $1-$ the normalized MSE or sMAPE for the methods on the weather and (subset of) M4 datasets, respectively.
 For classification and anomaly detection, we report the average accuracy and VUS-ROC of the methods across all the datasets.